### v1

# A universal scheme to self-test any quantum state and measurement

Abstract
The emergence of quantum devices has raised a significant issue: how to certify the quantum properties of a device without placing trust in it. To characterise quantum states and measurements in a device-independent way, up to some degree of freedom, we can make use of a technique known as self-testing. While schemes have been proposed to self-test all pure multipartite entangled states and real local rank-one projective measurements, little has been done to certify mixed entangled states, composite or non-projective measurements. By employing the framework of quantum networks, we propose a scheme that can be used to self-test any quantum state, projective measurement and rank-one extremal non-projective measurements. The quantum network considered in this work is the simple star network, which is implementable using current technologies. For our purposes, we also construct a family of Bell inequalities that can be used to self-test the two-dimensional tomographically complete set of measurements with an arbitrary number of parties.
Introduction.—Demonstrating the strength of quantum systems over classical ones in tasks like computation and cryptography has been one of the most important advancements in quantum information theory. These quantum protocols are more efficient and secure when compared to the classical ones. To implement these protocols, it is thus necessary to certify the quantumness of a given device. The traditional methods of certifying a quantum state or a measurement rely on trusting the devices involved in the scheme, which are referred to as device-dependent schemes. For instance, any quantum state can be device-dependently verified if one has complete knowledge of the measurement apparatuses. However, from a practical perspective, it is unreasonable to completely trust the device. Therefore, it is highly desirable to devise certification schemes where the assumptions made on the devices are minimal, also referred to as device-independent (DI) certification schemes. DI schemes allow one to verify the non-trivial quantum properties of an unknown device by only observing the statistical data it generates. The essential resources for any DI scheme are Bell nonlocal correlations that cannot be reproduced via any local (or simply classical) hidden variable models [1, 2, 3]. For instance, observation of the violation of a Bell inequality is a DI certification of the presence of entanglement inside the device.
Interestingly, the observation of Bell nonlocality in the data produced by a given device can provide a lot more information about the inner workings of the device than just the presence of entanglement. The strongest and most complete form of Bell nonlocality-based certification, known as self-testing [4, 5], allows for almost full characterization of the underlying quantum state and measurements performed on it. For this reason, in recent years, a plethora of schemes have been proposed to self-test pure quantum states or quantum measurements (see, e.g., Refs. [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]). In particular, Refs. [13] and [19] propose schemes for self-testing any bipartite or multipartite pure entangled states, where the second method is based on the quantum networks scenario. The only scheme to self-test a mixed entangled state, in particular, a bound entangled state, was only recently proposed in Ref. [20] (see nevertheless Refs. [21, 22]). As for quantum measurements, apart from a bunch of schemes allowing to certify various examples of quantum measurements [16, 14, 23, 17, 24], including the composite ones [25, 26, 27, 20], a general method to certify any real local rank-one projective measurement has only very recently been designed in Ref. [24]. Despite this progress, there exist no single unified scheme that allows for certification of quantum measurements and quantum states (pure or mixed). Moreover, no general scheme for quantum networks-based certification of composite measurements has been proposed.
The main aim of this work is to address the above challenges and provide a universal scheme that can be used to self-test any projective or rank-one extremal generalized quantum measurements (POVM) acting on arbitrary finite-dimensional Hilbert space as well as for any quantum state. For our purposes, we utilise the framework of quantum networks consisting of an arbitrary number of parties and sources (as depicted in Fig. 1 and explained later in the text). We first propose a Bell inequality that can be used to self-test the two-dimensional tomographically complete set of measurements. Based on this, we are then able to certify any projective or rank-one extremal generalized quantum measurement and eventually any quantum state. It is worth mentioning that, unlike the standard self-testing statements, our scheme requires additional causality constraints that are natural in quantum networks, in particular, that the sources are statistically independent. Our scheme is simple and can be implemented using the current infrastructure of quantum networks [28, 29, 30].
Before proceeding toward the main results, let us first introduce the scenario and relevant notions required throughout this work.
Preliminaries.—The quantum network scenario under scrutiny is composed of an external group of $N$ Alices, denoted $A_{i}$ $(i=1,\ldots,N)$, and a central party $E$ called Eve. All the $N+1$ parties are spatially separated. Correspondingly, $N$ independent sources distribute bipartite quantum states, in general, entangled states among the parties, as depicted in Fig. 1. These states are denoted by $\rho_{A_{i}E_{i}}$, where the subsystems $A_{i}$ will be measured by the external parties, whilst the other subsystems $E_{i}$ go to Eve. On their shares of the joint state $\rho_{AE}=\bigotimes_{i=1}^{N}\rho_{A_{i}E_{i}}$, where we denoted $A\coloneqq A_{1}\ldots A_{N}$ and $E\coloneqq E_{1}\ldots E_{N}$, each external party $A_{i}$ performs one of three available measurements, each with two outcomes; the measurement choices and outcomes of party $A_{i}$ are denoted $x_{i}=0,1,2$ and $a_{i}=0,1$, respectively. At the same time, the central party Eve freely chooses to perform one of two measurements where the first one has $2^{N}$ outcomes, whereas the second one has $K\leq 2^{N}$ outcomes. We denote Eve’s measurement choices by $e=0,1$ and outcomes by $l=0,\ldots,2^{N}-1/K$. Let us add here that the first Eve’s measurement is used to certify the external parties’ measurements as well as the states prepared by the sources whereas the second measurement is the one to be certified. Importantly, the parties cannot communicate classically during the experiment.
After running many rounds of the above measurements, all the parties can bring their data together to reconstruct the set of probability distributions $\vec{p}=\{p(\mathbf{a}l|\mathbf{x}e)\}$ which describe the correlations that are produced between all the observers, where each $p(\mathbf{a}l|\mathbf{x}e)$ is the probability of obtaining outcomes $a_{1}\ldots a_{N}=:\mathbf{a}$ by all the Alices and $l$ by Eve after they perform the measurements labelled by $x_{1}\ldots x_{N}=:\mathbf{x}$ and $e$, respectively. From Born’s rule, we can write
$p(\mathbf{a}l|\mathbf{x}e)=\mathrm{Tr}\left(\rho_{AE}\bigotimes_{i=1}^{N}M_{a_{i}|x_{i}}\otimes R_{l|e}\right),$ (1)
where $M_{i,x_{i}}=\{M_{a_{i}|x_{i}}\}$, for every $i$, and $E_{e}=\{R_{l|e}\}$ are the measurements performed by parties $A_{i}$ and $E$. The measurement elements $M_{a_{i}|x_{i}}$ and $R_{l|e}$ are positive semi-definite and sum up to the identity on the respective Hilbert space for every measurement choice $x_{i}$ or $e$ of all $N+1$ parties.
Figure 1: Depiction of the quantum network scenario. It consists of $N+1$ parties, namely, $A_{i}$ $(i=1,\ldots,N)$, and $E$, and $N$ independent sources distributing bipartite quantum states $\rho_{A_{i}E_{i}}$ among the parties as shown in the figure. The central party $E$ shares quantum states with each one of the other external parties $A_{i}$. While each $A_{i}$ has three inputs and two outcomes, $E$ has two inputs, The first Eve’s measurement has $12^{N}-$ 1 outcomes, whereas the second one has $K$ outcomes.
To better understand the observed correlations, expressing them in terms of the expected values of external observables can be helpful. These values are defined as
$\langle A_{1,x_{1}}\ldots A_{N,x_{N}}R_{l|e}\rangle=\sum_{a_{1}\ldots a_{N}=0}^{1}(-1)^{\sum_{i=1}^{N}a_{i}}p(\mathbf{a}l|\mathbf{x}e).$ (2)
It is important to note that by using Eq. (1), these expectation values can be expressed as $\langle A_{1,x_{1}}\ldots A_{N,x_{N}}R_{l|e}\rangle=\mathrm{Tr}[(\bigotimes_{i=1}^{N}A_{i,x_{i}})\otimes R_{l|e}\rho_{AE}]$, where $A_{i,x_{i}}$ are quantum operators defined via the measurement elements as $A_{i,x_{i}}=M_{0|x_{i}}-M_{1|x_{i}}$ for every $x_{i}$ and $i$. Notice that in a particular case when $\{M_{0|x_{i}},M_{1|x_{i}}\}$ is a projective measurement, the corresponding $A_{x_{i}}$ becomes a standard hermitian observable. For our convenience, in the above representation Eve’s measurements remain represented in terms of the measurement elements $R_{l|e}$ instead of observables.
Self-testing.—Let us introduce the idea of self-testing by referring back to the scenario shown in Fig. 1. First, we assume that the measurements conducted by the parties and the states $\rho_{A_{i}E_{i}}\in\mathcal{L}(\mathcal{H}_{A_{i}}\otimes\mathcal{H}_{E_{i}})$ for $i=1,\ldots,N$ are unknown. The only knowledge that Alices and Eve have about the whole system is encoded in the observed correlations $\vec{p}$. It is worth pointing out already now that since the dimensions of the local Hilbert spaces $\mathcal{H}_{A_{i}}$ and $\mathcal{H}_{E_{i}}$ are unspecified, we can utilise the standard dilation argument and presume that the shared states are pure, i.e., $\rho_{A_{i}E_{i}}=|\psi_{A_{i}E_{i}}\rangle\!\langle\psi_{A_{i}E_{i}}|$. On the same ground we can also assume that the measurements of the external parties are projective. However, this assumption can be dropped by employing a more general sum of squares decomposition of the Bell operators corresponding to the inequalities (7) which take into account that the measurements are not projective.
Let us then consider another, reference experiment giving rise to the same correlations $\vec{p}$ which is performed on known quantum states shared by the parties $|\psi^{\prime}_{A_{i}E_{i}}\rangle\in\mathcal{H}_{A_{i}^{\prime}}\otimes\mathcal{H}_{E_{i}^{\prime}}$ and with known measurements $A_{i,x_{i}}^{\prime}$ and $E^{\prime}_{e}$ of Alices and Eve, respectively, where $\mathcal{H}_{A_{i}^{\prime}}$ and $\mathcal{H}_{E_{i}^{\prime}}$ are some Hilbert spaces of known dimension. The task of self-testing is to deduce from the observed $\vec{p}$ that the actual experiment is equivalent to the reference one in the following sense: (i) the local Hilbert spaces admit the product form $\mathcal{H}_{A_{i}}=\mathcal{H}_{A_{i}^{\prime}}\otimes\mathcal{H}_{A_{i}^{\prime\prime}}$ and $\mathcal{H}_{E_{i}}=\mathcal{H}_{E_{i}^{\prime}}\otimes\mathcal{H}_{E_{i}^{\prime\prime}}$ for some auxiliary Hilbert spaces $\mathcal{H}_{A_{i}^{\prime\prime}}$ and $\mathcal{H}_{E_{i}^{\prime\prime}}$; and (ii) there are local unitary operations $U_{A_{i}}:\mathcal{H}_{A_{i}}\to\mathcal{H}_{A_{i}^{\prime}}\otimes\mathcal{H}_{A_{i}^{\prime\prime}}$ and $U_{E_{i}}:\mathcal{H}_{E_{i}}\to\mathcal{H}_{E_{i}^{\prime}}\otimes\mathcal{H}_{E_{i}^{\prime\prime}}$ such that
$(U_{A_{i}}\otimes U_{E_{i}})|\psi_{A_{i}E_{i}}\rangle=|\psi_{A_{i}^{\prime}E_{i}^{\prime}}\rangle\otimes|\xi_{A_{i}^{\prime\prime}E_{i}^{\prime\prime}}\rangle,$ (3)
where $|\xi_{A_{i}^{\prime\prime}E_{i}^{\prime\prime}}\rangle\in\mathcal{H}_{A_{i}^{\prime\prime}}\otimes\mathcal{H}_{E_{i}^{\prime\prime}}$ is some auxiliary quantum state, and
$U_{A_{i}}\,A_{i,x_{i}}\,U_{A_{i}}^{\dagger}=A_{i,x_{i}}^{\prime}\otimes\mathbbm{1}_{A_{i}^{\prime\prime}},\quad U_{E}\,R_{l|e}\,U_{E}^{\dagger}=R^{\prime}_{l|e}\otimes\mathbbm{1}_{E^{\prime\prime}},$ (4)
where $\mathbbm{1}_{E^{\prime\prime}}$ is the identity acting on the Eve’s auxiliary system $\mathcal{H}_{E^{\prime\prime}}$ and $U_{E}=\bigotimes_{i=1}^{N}U_{E_{i}}$ and we use the notation $E^{\prime\prime}=E_{1}^{\prime\prime}\ldots E_{N}^{\prime\prime}$.
If the above conditions (i) and (ii) are met one says that the reference state and measurements are self-tested in the actual experiment from the observed correlations. What is more, for any outcome $l$ of Eve’s measurement, the post-measurement states held by the external parties $A_{i}$,
$\tilde{\rho}^{l|e}_{A}=\frac{1}{\overline{P}(l|e)}\mathrm{Tr}_{E}\left[\left(\mathbbm{1}_{A}\otimes R_{l|e}\right)\bigotimes_{i=1}^{N}|\psi_{A_{i}E_{i}}\rangle\!\langle\psi_{A_{i}E_{i}}|\right],$ (5)
where $\overline{P}(l|e)$ is the probability that Eve observes the outcome $l$ after performing the $e-$th measurement satisfy
$U_{A}\,\rho^{l|e}_{A}\,U_{A}^{\dagger}=\tilde{\rho}^{l|e}_{A^{\prime}}\otimes\varrho_{A^{\prime\prime}}^{l|e},$ (6)
where $\varrho_{A^{\prime\prime}}^{l|e}$ are some auxiliary states and $U_{A}=U_{A_{1}}\otimes\ldots\otimes U_{A_{N}}$ where $U_{A_{i}}$ are the same unitary transformations as those in Eqs. (3) and (4), then we say that the post-measurement states are self-tested too. It should be noted here that the above definition of self-testing the sources (3), measurements (4), and the post-measurement states (6) holds up to the complex conjugation of the reference states and measurements. The reason is that the correlations $\vec{p}$ are invariant under complex conjugation of quantum states and measurements.
Let us finally mention that throughout this work, we assume that the local states of $|\psi_{A_{i}E_{i}}\rangle$ are full-rank as the local measurements can only be certified on the supports of the local density matrices.
Main results.—We are now ready to present our scheme for self-testing of composite measurements and quantum states. It is divided into three major parts. The first one involves self-testing of the two-dimensional tomographically complete set of Pauli measurements in the measurements of the external parties and the two-qubit Bell states in the states generated by all the sources. The second one involves self-testing any quantum measurement performed by the central party using the states and external parties’ measurements as certified in the above first part of the scheme. Lastly, the third part is concerned with applying the first two parts for self-testing the post-measurement state at the external parties’ laboratories. As a result, our schemes enable self-testing any quantum state (pure or mixed) distributed between the external parties by Eve’s measurements, while also facilitating the self-testing of quantum measurements performed by the central party.
Part 1. To self-test the tomographically complete set of measurements in the external parties’ measurement devices and the two-qubit singlets in the states generated by the sources, we introduce the following class of $2^{N}$ Bell inequalities that are suitable modifications of those introduced in Refs. [31, 20]:
$\mathcal{I}_{l}=(-1)^{l_{1}}\left[(N-1)\left\langle\tilde{A}_{1,1}\prod_{i=2}^{N}A_{i,1}\right\rangle+\sum_{i=2}^{N}(-1)^{l_{i}}\Biggl{\langle}\tilde{A}_{1,0}A_{i,0}\Biggl{\rangle}-(-1)^{l_{1}}\sum_{i=2}^{N}(-1)^{l_{i}}\left\langle A_{1,2}A_{i,2}\prod_{\begin{subarray}{c}j=2\\ j\neq i\end{subarray}}^{N}A_{j,1}\right\rangle\right]\leqslant\beta_{C},$ (7)
where $l\equiv l_{1}\ldots l_{N}$ such that $l_{1},l_{2},\ldots,l_{N}=0,1$ is the binary representation of the outcome $l$, and
$\tilde{A}_{1,0}=\frac{A_{1,0}-A_{1,1}}{\sqrt{2}},\qquad\tilde{A}_{1,1}=\frac{A_{1,0}+A_{1,1}}{\sqrt{2}}.$ (8)
Here, $A_{i,j}$ for $i=1,2,\ldots,N$ and $j=0,1,2$ are the observables measured by the external Alices.
Let us now briefly discuss the main properties of these Bell inequalities. First, it is direct to observe that their maximal values attainable using local deterministic strategies $\beta_{C}$, often referred to as the classical bound, are the same for any $l$ and amount to $\beta_{C}=(\sqrt{2}+1)(N-1)$ (see Fact 1 of [32] for a proof). Second, the maximal quantum values of $\mathcal{I}_{l}$, referred to as the quantum or Tsirelson’s bounds, equal $3(N-1)$ and are achieved by the following observables of the external parties
$A_{1,0}$ $=$ $\frac{X+Z}{\sqrt{2}},\qquad A_{1,1}=\frac{X-Z}{\sqrt{2}},\qquad A_{1,2}=Y,$
$A_{i,0}$ $=$ $Z,\qquad A_{i,1}=X,\qquad A_{i,2}=Y$ (9)
with $i=2,\ldots,N$ and the GHZ-like states
$|\phi_{l}\rangle=\frac{1}{\sqrt{2}}(|l_{1}\ldots l_{N}\rangle+(-1)^{l_{1}}|\overline{l}_{1}\ldots\overline{l}_{N}\rangle),$ (10)
where $l_{1},l_{2},\ldots,l_{N}=0,1$ and $\overline{l}_{i}=1-l_{i}$ (for a proof see Fact 2 in [32]). Notice that, the measurements in Eq. (A universal scheme to self-test any quantum state and measurement) form a two-dimensional tomographically complete set of measurements for every party. We now comprise the result in the following theorem:
Theorem 1.
Consider the scenario depicted in Fig. 1. If the Bell inequalities $\mathcal{I}_{l}$ (7) for any $l$ are maximally violated when Eve chooses the input $e=0$ with each outcome occurring with probability $\overline{P}(l|0)=1/2^{N}$, then the measurements of the external parties are equivalent, according to Eq. (4), to the reference measurements in (A universal scheme to self-test any quantum state and measurement), whereas the states $|\psi_{A_{i}E_{i}}\rangle$ are equivalent in the sense of Eq. (3) to the two-qubit maximally entangled state $|\phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$.
Sketch of the proof. Here we provide a sketch of the proof, deferring its full version to Appendix A. First, Eve performs the measurement corresponding to the input $e=0$ and observes an outcome $l$, in which way she prepares a post-measurement state shared by the external parties. Now, observation of maximal violation of the Bell inequalities $\mathcal{I}_{l}$ by the external parties allows one to self-test the post-measurement state as well as the external parties’ measurements. Then, the additional condition that the probabilities of obtaining the oucome $l$ by Eve when measuring $E_{0}$ is $1/2^{N}$ enables self-testing the states generated by the sources. This in turn also allows us to certify that Eve’s measurement $e=0$ is the Greenberger-Horne-Zeilinger (GHZ) basis $\{|\phi_{l}\rangle\!\langle\phi_{l}|\}$ where $|\phi_{l}\rangle$ are given in (10).
Part 2. Now, that the measurements performed by the external parties as well as the states generated by the sources are self-tested, we can show how the quantum network can be used to self-test any projective quantum measurement or rank-one extremal POVM in the second measurement $E_{1}$ performed by the central party (see Fig. 2). To this end, let us first consider the case of projective measurements and introduce a $K$-outcome reference measurement $E^{\prime}_{1}=\{R^{\prime}_{l|1}\}$ defined on $\mathcal{H}_{N}=\left(\mathbbm{C}^{2}\right)^{\otimes N}$. While we assume this measurement to be projective, we do not impose any restrictions on the rank $r_{l}$ of the measurement elements $R_{l|1}^{\prime}$.
Now, in order to certify the reference measurement $E_{1}^{\prime}$ in the actual one $E_{1}=\{R_{l|1}\}$, we require that the assumptions of Theorem 1 are satisfied and additionally that for any outcome $l$, the following condition
$\sum_{i_{1},\ldots,i_{N}=0}^{3}f_{i_{1},\ldots,i_{N}}^{l}\left\langle\tilde{A}_{1,i_{1}}\otimes\bigotimes_{k=2}^{N}A_{k,i_{k}}\otimes(R_{l|1})_{E}\right\rangle=\frac{r_{l}}{2^{N}},$ (11)
is satisfied by the correlations $\vec{p}$ observed in the network, where $f_{i_{1},\ldots,i_{N}}^{l}$ are real coefficients that come from expanding $R^{\prime}_{l|1}$ in the $N$-fold tensor products of the qubit Pauli matrices [See Appendix B of [32] for details]. Moreover, $\tilde{A}_{1,0}$ and $\tilde{A}_{1,1}$ are defined in (8), whereas $\tilde{A}_{1,2}=A_{1,2}$ and $\tilde{A}_{1,3}=A_{k,3}=\mathbbm{1}$ for any $k$. Notice that the above conditions are met in the ideal situation in which the sources generate the two-qubit maximally entangled state, the external parties measure the observables (19), and, finally, Eve’s measurement $E_{1}$ is such that $R_{l|1}=(R^{\prime}_{l|1})^{*}$. Consequently, we find the following result.
Figure 2: Self-testing any quantum measurement. First, we have self-tested all the measurements performed by each external party $A_{i}$ as (A universal scheme to self-test any quantum state and measurement), as well as, the sources as the maximally entangled two-qubit state $|\phi_{A_{i}E_{i}}^{+}\rangle$. Thus, Eve’s first measurement is certified to the one in the GHZ basis (10). These are the ingredients sufficient to certify Eve’s arbitrary second measurement.
Theorem 2.
Consider the scenario depicted in Fig. 2 with states generated by the sources and measurements of the external parties being self-tested as in Theorem 1. If the observed correlations $\vec{p}$ additionally satisfy the conditions (11) for Eve’s measurement $E_{1}=\{R_{l|1}\}$, then this measurement is self-tested to be the reference projective measurement $E^{\prime}_{1}$, according to the definition (4).
The proof can be found in Appendix B of [32]. Importantly, the above statement can be extended to any projective measurement defined on a $D$-dimensional Hilbert space with arbitrary $D$. In fact, one can always embed the measurement $\{R^{\prime}_{l|1}\}$ into $\mathcal{H}_{N}=(\mathbbm{C}^{2})^{\otimes N}$ where $N$ is the smallest natural number such that $D<2^{N}$. Then, in order to make the measurement fully supported on $\mathcal{H}_{N}$ one completes it with an additional projective measurement element $M^{\perp}$ such that $\sum_{l}R^{\prime}_{l|1}+M^{\perp}=\mathbbm{1}_{2^{N}}$. This gives rise to an $(K+1)$-outcome projective measurement to which Theorem 2 applies.
Interestingly, in a similar way one can also certify an arbitrary extremal rank-one POVM’s. For this purpose, let us assume that Eve’s reference measurement $E_{1}^{\prime}=\{R^{\prime}_{l|1}\}$ is now some $K$-outcome extremal POVM where the measurement elements are given by $R^{\prime}_{l|1}=\lambda_{l}|\phi_{l}\rangle\!\langle\phi_{l}|$ for some $0\leq\lambda_{l}\leq 1$ and pure $N$-qubit pure states $|\phi_{l}\rangle$ such that $\sum_{l}\lambda_{l}|\phi_{l}\rangle\!\langle\phi_{l}|=\mathbbm{1}$.
Now, to self-test this reference measurement in $E_{1}$ the parties check whether for all outcomes $l=1,\ldots,K$ and sequences $i_{1},\ldots,i_{N}$ with $i_{j}=0,1,2,3$, the correlations $\vec{p}$ satisfy the following conditions,
$\left\langle\tilde{A}_{1,i_{1}}\otimes\bigotimes_{k=2}^{N}A_{k,i_{k}}\otimes(R_{l|1})_{E}\right\rangle$ $=$ $f^{l}_{i_{1},\ldots,i_{N}},$ (12)
where $f^{l}_{i_{1},\ldots,i_{N}}$ are real numbers coming from the decompositions of the measurement operators $R^{\prime}_{l|1}$ in the $N$-fold tensor products of the qubit Pauli matrices. We can thus state our next result.
Theorem 3.
Consider the scenario depicted in Fig. 2 with states generated by the sources and measurements of the external parties being self-tested as in Theorem 1. If the observed correlations $\vec{p}$ additionally satisfy the conditions (12) for Eve’s measurement $E_{1}=\{R_{l|1}\}$, then this measurement is self-tested to be the reference rank-one extremal POVM $E^{\prime}_{1}$, according to the definition (4).
It should be noticed that similarly to the previous case, a rank-one extremal POVM $E_{1}^{\prime}$ defined on $\mathbbm{C}^{D}$ for some $D$ can be embedded in $\mathcal{H}_{N}$ with $N$ being the minimal number such that $D<2^{N}$. To this end, one completes the measurement with additional $2^{N}-D$ rank-one projections $J_{i}$ so that the resulting measurement $\{R^{\prime}_{1|1},\ldots,R^{\prime}_{K|1},J_{1},\ldots,J_{2^{N}-D}\}$ is also extremal (see Appendix C for more details).
We thus obtain a general way of certifying any projective or rank-one extremal generalized measurement in the quantum networks scenario. Our scheme is based on violation of a class of simple Bell inequalities (7) and some additional conditions that the observed correlations $\vec{p}$ must satisfy: (11) for projective measurements or (12) for rank-one extremal POVMs. It is however worth noticing that the generality of our scheme comes at the price of the complexity that grows fast with the dimension of the measurement under consideration and thus with the number of involved external parties.
Part 3. Finally, we are going to show now that using the above setup one can self-test any quantum state. Consider again the network scenario depicted in Fig. 2. It is clear that Eve could remotely prepare different quantum states with the external parties by performing the concerned quantum measurement on the maximally entangled states generated by the sources. Thus, without loss of generality, the sources along with Eve can be considered a new preparation device. For instance, consider that Eve performs a projective measurement $E_{1}=\{|e_{1}\rangle\!\langle e_{1}|,\ldots,|e_{K}\rangle\!\langle e_{K}|\}$. Then the post-measurement state shared by the external parties when she obtains an outcome $l$ is $|e_{l}^{*}\rangle\!\langle e_{l}^{*}|$ where *** represents the complex conjugate. As a corollary to the previous statements, we can state the following fact.
Corollary 1.
Consider the scenario depicted in Fig. 2. If the states generated by the sources along with the measurements performed by the external parties and Eve maximally violate the Bell inequalities (7) and satisfy the statistics (11) or (12) when Eve’s reference measurements are projective or rank-one extremal POVM respectively, then the post-measurement states with the external parties can be self-tested according to (6).
The proof of this statement can be found in Appendix D of [32]. It is important to emphasise here that, unlike the standard self-testing protocols, our scheme is not restricted to just pure entangled states. Rather, separable states and mixed states of any local dimension can also be self-tested with the external parties, according to (3).
It is simple to observe from the above Corollary that any pure state $|\psi\rangle$ can be self-tested with our scheme by appropriately choosing Eve’s projective measurement such that one of its measurement elements, say $R^{\prime}_{0|1}$ is exactly $R^{\prime}_{0|1}=|\psi\rangle\!\langle\psi|$.
Importantly, however, the scheme is not restricted to pure states only and can straightforwardly extended to mixed states. Consider a reference mixed state $\rho=\sum_{k}p_{k}|\psi_{k}\rangle\!\langle\psi_{k}|$ acting on $\mathbbm{C}^{d}$ which we aim to self-test. For this purpose, Eve needs to perform an extremal $3d-$outcome POVM in the Hilbert space $\mathbbm{C}^{2d}$ on her share of the joint state $|\psi_{AE}\rangle$ to create $\rho$ at the external parties’ labs with the aid of post-processing. To construct this POVM, we first define two sets of $d$ mutually orthogonal vectors $\{|\psi_{k}\rangle\}$ and $\{|\phi_{k}\rangle\}$ such that $\langle\psi_{k}|\phi_{k^{\prime}}\rangle=0$ for any $k,k^{\prime}$. One then considers a pair $\{|\psi_{k}\rangle,|\phi_{k}\rangle\}$ to construct the a three-outcome trine POVM [cf. Ref. [33]] as
$M_{k,1}$ $=$ $p_{k}|\psi_{k}\rangle\!\langle\psi_{k}|,\quad M_{k,2}=\alpha_{k}|\tau_{k,2}\rangle\!\langle\tau_{k,2}|,$
$M_{k,3}$ $=$ $\beta_{k}|\tau_{k,3}\rangle\!\langle\tau_{k,3}|,$ (13)
where for all $k$ the vectors $|\tau_{k,2}\rangle$ and $|\tau_{k,3}\rangle$ are certain linear combinations of the vectors $|\psi_{k}\rangle$ and $|\phi_{k}\rangle$ whose explicit forms are given in Eq. (160) in [32] and $\alpha_{k}=\beta_{k}=(2-p_{k})/2$. The desired extremal $3d$-outcome rank-one POVM is one which is composed of all the trine POVM’s, i.e., $E^{\prime}_{1}=\{M_{k,m}\}_{k,m}$.
Now, when Eve performs this measurement on her share of the joint state and obtain the outcomes $l\equiv(k,1)$ with probability $p_{k}/2^{N}$, the post-measurement (normalized) states of the external parties are given by $\tilde{\rho}_{(k,1)}=|\psi_{k}\rangle\!\langle\psi_{k}|$. These states are self-testable using Corollary 1. Then, the average state when Eve obtain the outcomes $(k,1)$ over all $k$ is thus certified to be
$U_{A}\,\rho_{A}\,U^{\dagger}_{A}=\frac{1}{\sum_{k}p(k,1)}\sum_{k}p(k,1)U_{A}\,\rho_{A}^{(k,1)}\,U_{A}^{\dagger}$
$=\left(\sum_{k}M_{(k,1)}\right)\otimes\tilde{\rho}_{A^{\prime\prime}}=\rho_{A^{\prime}}\otimes\tilde{\rho}_{A^{\prime\prime}},$ (14)
or its complex conjugate, which is exactly what we promised. Let us say that the external parties want to perform a task with the state $\rho$. Then, they perform their task with all the post-measurement states of Eve and after the task is done, just like Bell’s scenario Eve announces the outcomes. Now all the parties only consider the statistics when Eve obtains $(k,1)$ and discard the rest of the statistics.
Discussion. After completing our work, we have identified several follow-up issues. While we have provided a proof-of-principle scheme for self-testing any quantum state and extremal measurement, it is important to note that in actual experiments, the exact statistics mentioned may not be attainable, and instead, values that are close to them will be obtained. One crucial follow-up question is whether our scheme is resilient to experimental imperfections. Additionally, our current scheme requires highly entangled states and measurements, which can be costly to generate. Therefore, it would be interesting to explore the possibility of self-testing quantum states and measurements using partially entangled states and measurements. On the other hand, it would be interesting to generalize this scheme to any quantum measurement. Lastly, as we have utilised the framework of quantum networks, it would be valuable to investigate if our approach can be utilised for multiparty quantum cryptography or key distribution.
Acknowledgements.
This project was funded within the QuantERA II Programme (VERIqTAS project) that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733 and from the Polish National Science Center (project No 2021/03/Y/ST2/00175).
Supplementary Material: A universal scheme to self-test any quantum state and measurement
Appendix A Appendix A: Self-testing of N-GHZ bases, measurements of external parties, and the states prepared by the preparation devices
A.1 Bell inequalities
Let us consider the following $2^{N}$ Bell inequalities inspired from [31, 20]:
$\mathcal{I}_{l}=(-1)^{l_{1}}\left\langle(N-1)\tilde{A}_{1,1}\otimes\bigotimes_{i=2}^{N}A_{i,1}+\sum_{i=2}^{N}(-1)^{l_{i}}\tilde{A}_{1,0}\otimes A_{i,0}-(-1)^{l_{1}}\sum_{i=2}^{N}(-1)^{l_{i}}A_{1,2}\otimes A_{i,2}\otimes\bigotimes_{\begin{subarray}{c}j=2\\ j\neq i\end{subarray}}^{N}A_{j,1}\right\rangle\leqslant\beta_{C},$ (17)
where $l\equiv l_{1}l_{2}\ldots l_{N}$ with $l_{i}=0,1$ for each $i=1,\ldots,N$ is the binary representation of $l$, and
$\tilde{A}_{1,0}=\frac{A_{1,0}-A_{1,1}}{\sqrt{2}},\qquad\tilde{A}_{1,1}=\frac{A_{1,0}+A_{1,1}}{\sqrt{2}}.$ (18)
Here, $A_{i,j}$ for $i=1,2,\ldots,N$ and $j=0,1,2$ correspond to the observables of all the parties.
Fact 1.
The maximal classical value of the Bell expressions $\mathcal{I}_{l}$ for any $l$ is $\beta_{C}=(\sqrt{2}+1)(N-1)$.
Proof.
To compute the maximal classical value of $\mathcal{I}_{l}$ it is enough to optimize it over local deterministic strategies for which all the expectation values are products of the local ones, and, moreover, the local expectation values equal either $1$ or $-1$. Let us begin by noting that the maximal value of the last sum in Eq. (17) over such strategies is $N-1$ and it is independent of the choice of the maximal values of the first two term. To determine the latter we can to consider two cases: $A_{1,0}=A_{1,1}$ and $A_{1,0}\neq A_{1,1}$. In the first one, the maximal value of the first term in Eq. (17) is $\sqrt{2}(N-1)$, whereas the second term vanishes. In the second case, the first term in $\mathcal{I}_{l}$ vanishes, whereas the maximal value of the second term is again $\sqrt{2}(N-1)$. Thus, $\beta_{C}=(\sqrt{2}+1)(N-1)$. ∎
Fact 2.
The maximal quantum value of the Bell expression $\mathcal{I}_{l}$ is $3(N-1)$ and it is achieved by the following observables
$A_{1,0}=\frac{X+Z}{\sqrt{2}},\quad A_{1,1}=\frac{X-Z}{\sqrt{2}},\quad A_{1,2}=Y,\quad A_{i,0}=Z,\quad A_{i,1}=X,\quad A_{i,2}=Y\quad(i=2,\ldots,N)$ (19)
as well as the GHZ-like states
$|\phi_{l}\rangle=\frac{1}{\sqrt{2}}(|l_{1}\ldots l_{N}\rangle+(-1)^{l_{1}}|\overline{l}_{1}\ldots\overline{l}_{N}\rangle),$ (20)
where $l\equiv l_{1}l_{2}\ldots l_{N}$ with $l_{1},l_{2},\ldots,l_{N}=0,1$ and $\overline{l}_{i}=1-l_{i}$.
Proof.
The following proof is inspired by Refs. [31, 20]. Let us first associate a Bell operator to each of the Bell expressions $\mathcal{I}_{l}$ (17) of the following form
[…]
where
$P_{l_{1}}=(-1)^{l_{1}}\tilde{A}_{1,1}-\bigotimes_{i=2}^{N}A_{i,1},$ (23a)
$R_{i,l_{1},l_{i}}=(-1)^{l_{1}+l_{i}}\tilde{A}_{1,0}-A_{i,0},$ (23b)
$Q_{i,l_{i}}=(-1)^{l_{i}}A_{1,2}+A_{i,2}\otimes\bigotimes_{\begin{subarray}{c}j=2\\ j\neq i\end{subarray}}^{N}A_{j,1}.$ (23c)
Notice that by expanding the terms on the right-hand side of the above decomposition (22) one obtains
[…]
where to arrive at the last line we used the fact that $\tilde{A}_{i,j}^{2}\leqslant\mathbbm{1}$. Thus, one can conclude from (26) that $3(N-1)$ is an upper bound on $\mathcal{I}_{l}$ when using quantum theory. One can easily verify that the Bell inequalities (17) attain the value $3(N-1)$ when the states shared among the parties are the GHZ-like state $|\phi_{l}\rangle$ (20) and the observables given in Eq. (19). Thus, the Tsirelson’s bound of the Bell inequalities (17) for any $l$ is $3(N-1)$. ∎
An important observation from the above SOS decomposition is that when the maximal violation of the above Bell inequalities is attained, the right-hand side of Eq. (22) is 00. Thus, any state $|\psi\rangle$ and observables $A_{i,j}$ that attain the maximal violation of the Bell inequalities (17) for any $l$ is given by
$(-1)^{l_{1}}\tilde{A}_{1,1}|\psi\rangle=\bigotimes_{i=2}^{N}A_{i,1}|\psi\rangle,$ (28a)
$(-1)^{l_{1}+l_{i}}\tilde{A}_{1,0}|\psi\rangle=A_{i,0}|\psi\rangle\qquad(i=2,\ldots,N),$ (28b)
$-(-1)^{l_{i}}A_{1,2}|\psi\rangle=A_{i,2}\otimes\bigotimes_{\begin{subarray}{c}j=2\\ j\neq i\end{subarray}}^{N}A_{j,1}|\psi\rangle\qquad(i=2,\ldots,N).$ (28c)
Also, one can conclude from Eq. (26) that the observables that attain the maximum violation of the Bell inequalities (17) must be unitary, that is, $A_{i,j}^{2}=\mathbbm{1}$. These relations will be particularly useful for self-testing.
A.2 Self-testing
Theorem 1.
Assume that the Bell inequalities (17) for any $l$ are maximally violated and each outcome of the central party occurs with probability $\overline{P}(l|0)=1/2^{N}$ where $N$ denotes the number of external parties. The sources $P_{i}$ prepare the states $|\psi_{A_{i}E_{i}}\rangle\in\mathcal{H}_{A_{i}}\otimes\mathcal{H}_{E_{i}}$ for $i=1,2,\ldots,N$, the measurement with the central party is given by $\{R_{l|0}\}$ for $l=l_{1}l_{2}\ldots l_{N}$ such that $l_{i}=0,1$ which acts on $\bigotimes_{i=1}^{N}\mathcal{H}_{E_{i}}$, the local observables for each of the party is given by $A_{i,0}$, $A_{i,1}$, and $A_{i,2}$ for $i=1,2,\ldots,N$ which act on $\mathcal{H}_{A_{i}}$. Then,
  1. 1.
    The Hilbert spaces of all the parties decompose as $\mathcal{H}_{A_{i}}=\mathcal{H}_{A_{i}^{\prime}}\otimes\mathcal{H}_{A_{i}^{\prime\prime}}$ and $\mathcal{H}_{E_{i}}=\mathcal{H}_{E_{i}^{\prime}}\otimes\mathcal{H}_{E_{i}^{\prime\prime}}$, where $\mathcal{H}_{A_{i}^{\prime}}$ and $\mathcal{H}_{E_{i}^{\prime}}$ are qubit Hilbert spaces, whereas $\mathcal{H}_{A_{i}^{\prime\prime}}$ and $\mathcal{H}_{E_{i}^{\prime\prime}}$ are some finite-dimensional but unknown auxiliary Hilbert spaces.
  2. 2.
    There exist local unitary transformations $U_{A_{i}}:\mathcal{H}_{A_{i}}\rightarrow(\mathbb{C}^{2})_{A^{\prime}}\otimes\mathcal{H}_{A^{\prime\prime}_{i}}$ and $U_{E_{i}}:\mathcal{H}_{E_{i}}\rightarrow(\mathbb{C}^{2})_{A^{\prime}}\otimes\mathcal{H}_{E^{\prime\prime}_{i}}$ such that the states are given by
    $U_{A_{i}}\otimes U_{E_{i}}|\psi_{A_{i}E_{i}}\rangle=|\phi^{+}_{A_{i}^{\prime}E^{\prime}_{i}}\rangle\otimes|\xi_{A_{i}^{\prime\prime}E^{\prime\prime}_{i}}\rangle.$ (29)
  3. 3.
    The measurement of the central party is
    $U_{E}R_{l}U_{E}^{\dagger}=|\phi_{l}\rangle\!\langle\phi_{l}|_{E^{\prime}}\otimes\mathbbm{1}_{E^{\prime\prime}}\qquad\forall l,$ (30)
    where $U_{E}=U_{E_{1}}\otimes\ldots\otimes U_{E_{N}}$ and $|\phi_{l}\rangle$ are given in Eq. (20). The measurements of all the other parties are given by
    $U_{A_{1}}A_{1,0}\,U_{A_{1}}^{\dagger}$ $=$ $\left(\frac{X+Z}{\sqrt{2}}\right)_{A_{1}^{\prime}}\otimes\mathbbm{1}_{A_{1}^{\prime\prime}},\quad U_{A_{1}}A_{1,1}\,U_{A_{1}}^{\dagger}=\left(\frac{X-Z}{\sqrt{2}}\right)_{A_{1}^{\prime}}\otimes\mathbbm{1}_{A_{1}^{\prime\prime}},$
    $U_{A_{i}}A_{i,0}\,U_{A_{i}}^{\dagger}$ $=$ $Z_{A_{i}^{\prime}}\otimes\mathbbm{1}_{A_{i}^{\prime\prime}},\quad\quad\ \ \ \ \ U_{A_{i}}A_{i,1}\,U_{A_{i}}^{\dagger}=X_{A_{i}^{\prime}}\otimes\mathbbm{1}_{A_{i}^{\prime\prime}}\qquad(i=2,\ldots,N)$ (31)
    and,
    $U_{A_{i}}A_{i,2}\,U_{A_{i}}^{\dagger}$ $=$ ±$\pm Y_{A_{i}^{\prime}}\otimes\mathbbm{1}_{A_{i}^{\prime\prime}}\qquad\forall i.$ (32)
Proof.
Before proceeding, let us recall that the state shared among the external parties $A_{i}$ when the central party $E_{i}$ gets an outcome $l$ is given by
$\rho^{l}_{{A}}=\frac{1}{\overline{P}(l|0)}\mathrm{Tr}_{E}\left[\left(\mathbbm{1}_{A}\otimes R_{l}\right)\bigotimes_{i=1}^{N}|\psi_{A_{i}E_{i}}\rangle\!\langle\psi_{A_{i}E_{i}}|\right],$ (33)
where we denoted $A_{1}\ldots A_{N}\equiv A$ and $E_{1}\ldots E_{N}\equiv E$. Similar notation will be used for $A^{\prime}$, $A^{\prime\prime}$, $E^{\prime}$, and $E^{\prime\prime}$. Now, we first find the states $\rho^{l}_{{A}}$ for any $l$ and characterise the measurements $A_{i,j}$ for any $i$ and $j=0,1$. Then, using these derived states, we certify the states generated by the sources $|\psi_{A_{i}E_{i}}\rangle$ for all $i$. Then, using both of these results we certify the measurement $\{R_{l|0}\}$. For simplicity, we represent $R_{l|0}\equiv R_{l}$ in the below proof. Finally, using all the derived results, we characterise the observables $A_{i,2}$.
A.   Post-measurement states $\rho^{l}_{{A}}$
To characterize the post-measurement states $\rho^{l}_{{A}}$ based on violation of our Bell inequalities we need to consider the relations (28a) and (28b) obtained from the sum of squares decomposition of the corresponding Bell operators. First, let us consider a purification of the state $\rho^{l}_{{A}}$ which is obtained by adding an ancillary system $G$ such that
$\rho^{l}_{{A}}=\mathrm{Tr}_{G}\left(|\psi_{l}\rangle\!\langle\psi_{l}|_{{AG}}\right).$ (34)
In what follows, for simplicity, we drop the subscript $AG$ from the state $|\psi_{l}\rangle_{AG}$.
Now, for each outcome of the measurement performed by the central party, the corresponding Bell inequalities (17) are maximally violated by the post-measurement states $|\psi_{l}\rangle$. Thus, these states and the observables of every party must satisfy the relations (28a) and (28b), which we state as:
$(-1)^{l_{1}}\frac{A_{1,0}+A_{1,1}}{\sqrt{2}}|\psi_{l}\rangle$ $=\bigotimes_{i=2}^{N}A_{i,1}|\psi_{l}\rangle,$ (35a)
$(-1)^{l_{1}+l_{i}}\frac{A_{1,0}-A_{1,1}}{\sqrt{2}}|\psi_{l}\rangle$ $=A_{i,0}|\psi_{l}\rangle$ (35b)
for any $l=l_{1}l_{2}\ldots l_{N}$ and $i=2,3,\ldots,N$. Notice that any measurement can only be characterised on the support of the local states of $|\psi_{l}\rangle$. Thus, let us denote by $\Pi_{l}^{A_{1}}$ the projection onto the support of $\rho_{A_{1}}^{l}=\mathrm{Tr}_{A\setminus A_{1}}(\rho_{A}^{l})$ and apply it to the above equations to obtain:
$(-1)^{l_{1}}\frac{\mathbb{A}_{1,0}^{(l)}+\mathbb{A}_{1,1}^{(l)}}{\sqrt{2}}|\psi_{l}\rangle$ $=\bigotimes_{i=2}^{N}A_{i,1}|\psi_{l}\rangle,$ (36a)
$(-1)^{l_{1}+l_{i}}\frac{\mathbb{A}_{1,0}^{(l)}-\mathbb{A}_{1,1}^{(l)}}{\sqrt{2}}|\psi_{l}\rangle$ $=A_{i,0}|\psi_{l}\rangle\qquad(i=2,3,\ldots,N),$ (36b)
where we introduce the notation $\mathbb{A}_{1,j}^{(l)}=\Pi_{l}^{A_{1}}A_{1,j}\Pi_{l}^{A_{1}}$ for $j\in\{0,1\}$. Observe that $\Pi_{l}^{A_{1}}|\psi_{l}\rangle=|\psi_{l}\rangle$.
Let us now recall that to achieve the maximal violation of the Bell inequalities (17) the observables must be unitary, that is, $A_{i,j}^{2}=\mathbbm{1}$. Now, consider the relations (36a) and (36b), and multiply them by $\bigotimes_{i=2}^{N}A_{i,1}$ and $A_{i,0}$, respectively. By using the same relations on themselves again, we obtain that:
$\left(\mathbb{A}_{1,0}^{(l)}+\mathbb{A}_{1,1}^{(l)}\right)^{2}|\psi_{l}\rangle$ $=2|\psi_{l}\rangle;$ (37a)
$\left(\mathbb{A}_{1,0}^{(l)}-\mathbb{A}_{1,1}^{(l)}\right)^{2}|\psi_{l}\rangle$ $=2|\psi_{l}\rangle.$ (37b)
We can assume that the local states $\rho_{A_{1}}$ are full-rank and thus invertible. Therefore, if we take a partial trace over all the other subsystems except the first one, we obtain from (37a) and (37b) that
$\left(\mathbb{A}_{1,0}^{(l)}+\mathbb{A}_{1,1}^{(l)}\right)^{2}=2\Pi_{l}^{A_{1}}$ (38)
and
$\left(\mathbb{A}_{1,0}^{(l)}-\mathbb{A}_{1,1}^{(l)}\right)^{2}=2\Pi_{l}^{A_{1}},$ (39)
which results in
$\left\{\mathbb{A}_{1,0}^{(l)},\mathbb{A}_{1,1}^{(l)}\right\}=0.$ (40)
Also, because $\left(\mathbb{A}_{1,i}^{(l)}\right)^{2}\leqslant\Pi_{l}^{A_{1}}$, the above equations imply that $\left(\mathbb{A}_{1,i}^{(l)}\right)^{2}=\Pi_{l}^{A_{1}}$ for $i\in\{0,1\}$. This means that $\mathbb{A}_{1,i}^{(l)}$ are unitary on the support of $\rho_{A_{1}}^{l}$ and and a consequence, the observables $A_{1,i}$ can be represented as direct sums,
$A_{1,i}=\mathbb{A}_{1,i}^{(l)}\oplus F_{i}^{(l)},$ (41)
where $F_{i}^{(l)}$ act on the complement of $\mathrm{supp}(\rho_{A_{1}}^{l})$ in $\mathcal{H}_{A_{1}}$. This means that the anticommutation relations (40) can also be stated as
$\left\{\mathbb{A}_{1,0}^{(l)},\mathbb{A}_{1,1}^{(l)}\right\}=\Pi_{l}^{A_{1}}\left\{A_{1,0},A_{1,1}\right\}=0,$ (42)
which allows us to conclude that
$\sum_{l}\Pi_{l}^{A_{1}}\left\{A_{1,0},A_{1,1}\right\}=0.$ (43)
Let us now show that $\sum_{l}\Pi_{l}^{A_{1}}>0$, that is, the sum of all projections $\Pi_{l}^{A_{1}}$ is invertible. To see this, suppose that $\sum_{l}\Pi_{l}^{A_{1}}\ngtr 0$. That implies that exists $|\phi\rangle$ such that $\sum_{l}\Pi_{l}|\phi\rangle=0$. Thus, for every $l$, we must have $\Pi_{l}|\phi\rangle=0$, which means that $|\phi\rangle$ belongs to the kernel of $\rho_{A_{1}}^{l}$, i.e., $\rho_{A_{1}}^{l}|\phi\rangle=0$. By summing over $l$, we obtain $\rho_{A_{1}}|\phi\rangle=0$, which contradicts the assumption that the local states of $|\psi_{A_{i}E_{i}}\rangle$ are full rank. Therefore, we must have
$\sum_{l}\Pi_{l}^{A_{1}}>0,$ (44)
which implies that
$\{A_{1,0},A_{1,1}\}=0.$ (45)
As proven in e.g. Ref. [14], for a pair of unitary observables with eigenvalues ±$\pm 1$ that anti-commute, there exist a local unitary $V_{A_{1}}:\mathcal{H}_{A_{1}}\rightarrow\mathcal{H}_{A_{1}}$ such that
$V_{A_{1}}A_{1,0}V_{A_{1}}^{\dagger}$ $=$ $Z_{A_{1}^{\prime}}\otimes\mathbbm{1}_{A_{1}^{\prime\prime}},\qquad V_{A_{1}}A_{1,1}V_{A_{1}}^{\dagger}=X_{A_{1}^{\prime}}\otimes\mathbbm{1}_{A_{1}^{\prime\prime}}.$ (46)
Now, we can always apply a unitary $V^{\prime}:\mathcal{H}_{A_{1}^{\prime}}\rightarrow\mathcal{H}_{A_{1}^{\prime}}$ such that $V^{\prime}ZV^{\prime\dagger}=(X+Z)/\sqrt{2}$ and $V^{\prime}XV^{\prime\dagger}=(X-Z)/\sqrt{2}$. Thus, denoting $U_{A_{1}}=(V^{\prime}\otimes\mathbbm{1}_{A_{1}^{\prime\prime}})\,V_{A_{1}}$, we obtain that there exists a unitary operation $U_{A_{1}}:\mathcal{H}_{A1}\to(\mathbbm{C}^{2})_{A_{1}^{\prime}}\otimes\mathcal{H}_{A_{1}^{\prime\prime}}$ for which,
$U_{A_{1}}A_{1,0}\,U_{A_{1}}^{\dagger}$ $=$ $\frac{X+Z}{\sqrt{2}}\otimes\mathbbm{1}_{A_{1}^{\prime\prime}},\qquad U_{A_{1}}A_{1,1}\,U_{A_{1}}^{\dagger}=\frac{X-Z}{\sqrt{2}}\otimes\mathbbm{1}_{A_{1}^{\prime\prime}}.$ (47)
Now, let us derive the form of the observables $A_{i,0}$ and $A_{i,1}$ of all the other parties. For this purpose, let us first consider the $k$-th party and then use the derived observables (47) to rewrite the relations (35a) and (35b) as
$(-1)^{l_{1}}X_{A_{1}^{\prime}}|\psi^{\prime}_{l}\rangle=\mathbb{A}_{k,1}^{(l)}\otimes\bigotimes_{i=2\setminus\{k\}}^{N}A_{i,1}|\psi^{\prime}_{l}\rangle$ (48)
and,
$(-1)^{l_{1}+l_{k}}Z_{A_{1}^{\prime}}|\psi^{\prime}_{l}\rangle=\mathbb{A}_{k,0}^{(l)}|\psi^{\prime}_{l}\rangle,$ (49)
where $|\psi^{\prime}_{l}\rangle=U_{A_{1}}|\psi_{l}\rangle$, and $\mathbb{A}_{k,i}^{(l)}=\Pi_{l}^{A_{k}}A_{k,i}\,\Pi_{l}^{A_{k}}$ for $i\in\{0,1\}$, is the projected operator defined in analogous way as before. Now, multiplying (48) with $Z_{A_{1}^{\prime}}$ and then using (49) on the right-hand side of the obtained expression, we get
$(-1)^{l_{k}}(ZX)_{A_{1}^{\prime}}|\psi^{\prime}_{l}\rangle=\mathbb{A}_{k,1}^{(l)}\mathbb{A}_{k,0}^{(l)}\otimes\bigotimes_{i=2\setminus\{k\}}^{N}A_{i,1}|\psi^{\prime}_{l}\rangle.$ (50)
Now, multiplying (49) with $X_{A_{1}^{\prime}}$ and then using (48) on the right-hand side of the obtained expression, we get
$(-1)^{l_{k}}(XZ)_{A_{1}^{\prime}}|\psi^{\prime}_{l}\rangle=\mathbb{A}_{k,0}^{(l)}\mathbb{A}_{k,1}^{(l)}\otimes\bigotimes_{i=2\setminus\{k\}}^{N}A_{i,1}|\psi^{\prime}_{l}\rangle.$ (51)
Adding (50) and (51) and using the fact that the Pauli matrices satisfy $ZX+XZ=0$, we finally obtain that
$\left\{\mathbb{A}_{k,0}^{(l)},\mathbb{A}_{k,1}^{(l)}\right\}\otimes\bigotimes_{i=2\setminus\{k\}}^{N}A_{i,1}|\psi^{\prime}_{l}\rangle=0.$ (52)
Now, using the fact that the observables $A_{i,j}$ are unitary, we get that
$\left\{\mathbb{A}_{k,0}^{(l)},\mathbb{A}_{k,1}^{(l)}\right\}=0\qquad(k=2,\ldots,N).$ (53)
By applying $X_{A_{1}^{\prime}}$ and $Z_{A_{1}^{\prime}}$ to Eqs. (48) and (49), respectively, we obtain that $(\mathbb{A}_{k,i}^{(l)})^{2}=\Pi_{l}^{A_{k}}$. Analogously to the previous argument, we can conclude that
$\left\{A_{k,0},A_{k,1}\right\}=0\qquad(k=2,\ldots,N).$ (54)
Again, using the result presented in the appendix of [14], we observe that there exists unitaries $U_{A_{i}}:\mathcal{H}_{A_{i}}\rightarrow\mathcal{H}_{A_{i}}$ such that
$U_{A_{k}}A_{k,0}\,U_{A_{k}}^{\dagger}=Z_{A_{i}^{\prime}}\otimes\mathbbm{1}_{A_{i}^{\prime\prime}},\qquad U_{A_{k}}A_{k,1}\,U_{A_{k}}^{\dagger}=X_{A_{i}^{\prime}}\otimes\mathbbm{1}_{A_{i}^{\prime\prime}}\qquad(k=2,\ldots,N).$ (55)
Now, let us characterise the states $\rho^{l}_{{A}}$ for any $l$. For this purpose, we again consider the relations (28a) and (28b) and substitute into them the derived observables from (47) and (55). This gives us
$(-1)^{l_{1}}\bigotimes_{i=1}^{N}X_{A_{i}^{\prime}}|\tilde{\psi}_{l}\rangle=|\tilde{\psi}_{l}\rangle$ (56)
and,
$(-1)^{l_{1}+l_{i}}Z_{A_{1}^{\prime}}\otimes Z_{A_{i}^{\prime}}|\tilde{\psi}_{l}\rangle=|\tilde{\psi}_{l}\rangle\qquad(i=2,\ldots,N),$ (57)
where $|\tilde{\psi}_{l}\rangle=(U_{A_{1}}\otimes\ldots\otimes U_{A_{N}})|\psi_{l}\rangle$. As concluded above, the local Hilbert spaces $\mathcal{H}_{A_{i}}$ are even-dimensional, and therefore any state $|\tilde{\psi}_{l}\rangle$ belonging to the Hilbert space $\bigotimes_{i=1}^{N}\mathcal{H}_{A_{i}}\otimes\mathcal{H}_{G}$ can be written as
$|\tilde{\psi}_{l}\rangle=\sum_{i_{1},i_{2},\ldots i_{N}=0,1}|i_{1}i_{2}\ldots i_{N}\rangle_{A^{\prime}}|\phi^{l}_{i_{1}i_{2}\ldots i_{N}}\rangle_{A^{\prime\prime}G},$ (58)
where $|\phi^{l}_{i_{1}i_{2}\ldots i_{N}}\rangle_{A^{\prime\prime}G}$ are some unknown (in general unnormalized) states corresponding to the auxiliary degrees of freedom $A_{i}^{\prime\prime}$ and the system $G$ which is responsible for purification.
Putting the state (58) into the conditions (57), we obtain
$(-1)^{l_{1}+l_{k}}\sum_{i_{1},i_{2},\ldots i_{N}=0,1}(-1)^{i_{1}+i_{k}}|i_{1}i_{2}\ldots i_{N}\rangle|\phi^{l}_{i_{1}i_{2}\ldots i_{N}}\rangle=\sum_{i_{1},i_{2},\ldots i_{N}=0,1}|i_{1}i_{2}\ldots i_{N}\rangle|\phi^{l}_{i_{1}i_{2}\ldots i_{N}}\rangle,$ (59)
where $|\phi^{l}_{i_{1}i_{2}\ldots i_{N}}\rangle$ are in general unnormalised. Projecting the above formula on $\langle i_{1}i_{2}\ldots i_{N}|$, we get
$(-1)^{l_{1}+l_{k}}(-1)^{i_{1}+i_{k}}|\phi^{l}_{i_{1}i_{2}\ldots i_{N}}\rangle=|\phi^{l}_{i_{1}i_{2}\ldots i_{N}}\rangle\qquad(k=2,\ldots,N).$ (60)
From the above condition, we can conclude that $|\phi^{l}_{i_{1}i_{2}\ldots i_{N}}\rangle=0$ whenever $l_{1}+l_{k}+i_{1}+i_{k}$ mod $2=1$ for any $k=2,3,\ldots,N$. Thus, the state (58) that satisfies the condition (60) is given by
$|\tilde{\psi}_{l}\rangle=|l_{1}\ldots l_{N}\rangle|\phi_{l_{1}\ldots l_{N}}\rangle+|\overline{l}_{1}\ldots\overline{l}_{N}\rangle|\phi_{\overline{l}_{1}\ldots\overline{l}_{N}}\rangle,$ (61)
where $l_{i}=0,1$ for any $i=1,2,\ldots,N$ and $\overline{l}_{i}$ denotes the negation of the bit $l_{i}$. Now, putting this state (61) to the condition (56), we obtain that
$(-1)^{l_{1}}\left[|\overline{l}_{1}\ldots\overline{l}_{N}\rangle|\phi_{l_{1}\ldots l_{N}}\rangle+|l_{1}\ldots l_{N}\rangle|\phi_{\overline{l}_{1}\ldots\overline{l}_{N}}\rangle\right]=|l_{1}\ldots l_{N}\rangle|\phi_{l_{1}\ldots l_{N}}\rangle+|\overline{l}_{1}\ldots\overline{l}_{N}\rangle|\phi_{\overline{l}_{1}\ldots\overline{l}_{N}}\rangle.$
This implies that
$|\phi_{\overline{l}_{1}\ldots\overline{l}_{N}}\rangle=(-1)^{l_{1}}|\phi_{l_{1}\ldots l_{N}}\rangle,$ (63)
which after substituting into Eq. (61) allows us to conclude that the state $|\tilde{\psi}_{l}\rangle$ is given by
$|\tilde{\psi}_{l}\rangle=\frac{1}{\sqrt{2}}\left(|l_{1}l_{2}\ldots l_{N}\rangle+(-1)^{l_{1}}|\overline{l}_{1}\ldots\overline{l}_{N}\rangle\right)_{A^{\prime}}\otimes|\phi_{l}\rangle_{A^{\prime\prime}G},$ (64)
where we have also included the normalization factor. Tracing the $G$ subsystem out, we finally obtain
$\left(\bigotimes_{i=1}^{N}U_{A_{i}}\right)\rho^{l}_{{A}}\left(\bigotimes_{i=1}^{N}U_{A_{i}}^{\dagger}\right)=|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\tilde{\rho}^{l}_{A^{\prime\prime}}.$ (65)
where $\tilde{\rho}^{l}_{A^{\prime\prime}}$ denotes the auxiliary state acting on $\bigotimes_{i=1}^{N}\mathcal{H}_{A_{i}^{\prime\prime}}$.
Putting it back into Eq. (33) and taking the unitaries to the right-hand side, we see that
$|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\tilde{\rho}^{l}_{A^{\prime\prime}}=2^{N}\mathrm{Tr}_{E}\left[\left(\mathbbm{1}_{A}\otimes R_{l}\right)\bigotimes_{i=1}^{N}|\tilde{\psi}_{A_{i}E_{i}}\rangle\!\langle\tilde{\psi}_{A_{i}E_{i}}|\right],$ (66)
where $|\tilde{\psi}_{A_{i}E_{i}}\rangle=U_{A_{i}}|\psi_{A_{i}E_{i}}\rangle$ and we used the fact that $\overline{P}(l|0)=1/2^{N}$.
B.  States generated by the source $|\psi_{A_{i}E_{i}}\rangle$
Our aim now is to characterize the states produced by the sources $|\psi_{A_{i}E_{i}}\rangle$ and to this aim we follow the approach of Ref. [20]. First, with the aid of the Schmidt decomposition each state $|\tilde{\psi}_{A_{i}E_{i}}\rangle=U_{A_{i}}|\psi_{A_{i}E_{i}}\rangle$ can be expressed in the following way
$|\tilde{\psi}_{A_{i}E_{i}}\rangle=\sum_{j=0}^{d_{i}-1}\alpha_{j}|e_{j}\rangle_{A_{i}}|f_{j}\rangle_{E_{i}}$ (67)
such that $d_{i}$ denotes the dimension of the Hilbert space $\mathcal{H}_{A_{i}}$ and $\{|e_{j}\rangle_{A_{i}}\}$ and $\{|f_{j}\rangle_{E_{i}}\}$ denotes the local orthonormal bases of the subsystem $A_{i}$ and $E_{i}$, respectively. Notice that, as already proven, the dimension $d_{i}$ is even for any $i$ as the Hilbert space of any external party decomposes as $\mathcal{H}_{A_{i}}=\mathbbm{C}^{2}\otimes\mathcal{H}_{A_{i}^{\prime\prime}}$. Moreover, the Schmidt coefficients satisfy $\alpha_{j}>0$ and $\sum_{j}\alpha_{j}^{2}=1$.
Let us now consider a unitary operation $U_{E_{i}}:\mathcal{H}_{E_{i}}\rightarrow\mathcal{H}_{E_{i}}$ such that $U_{E_{i}}|f_{j}\rangle=|e_{j}^{*}\rangle_{E_{i}}$. After applying it to the state (67), one arrives at
$|\overline{\psi}_{A_{i}E_{i}}\rangle=U_{E_{i}}|\tilde{\psi}_{A_{i}E_{i}}\rangle=\sum_{j=0}^{d_{i}-1}\alpha_{j}|e_{j}\rangle_{A_{i}}|e_{j}^{*}\rangle_{E_{i}}$ (68)
Let us now define a matrix $P_{E_{i}}$ of rank $d_{i}$ as,
$P_{E_{i}}=\sum_{j=0}^{d_{i}-1}\alpha_{j}|e_{j}^{*}\rangle\!\langle e_{j}^{*}|,$ (69)
using which we rewrite the state (68) as
[…]
Here $|\phi^{+,d_{i}}_{A_{i}E_{i}}\rangle$ denotes the maximally entangled state of local dimension $d_{i}$, that is,
$|\phi^{+,d_{i}}_{A_{i}E_{i}}\rangle=\frac{1}{\sqrt{d_{i}}}\sum_{j=0}^{d_{i}-1}|e_{j}\rangle_{A_{i}}|e_{j}^{*}\rangle_{E_{i}}=\frac{1}{\sqrt{d_{i}}}\sum_{j=0}^{d_{i}-1}|j\rangle_{A_{i}}|j\rangle_{E_{i}}.$ (71)
Putting it back into the condition (66), we get that
$|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\tilde{\rho}^{l}_{A^{\prime\prime}}=2^{N}\mathrm{Tr}_{E}\left[\left(\mathbbm{1}_{A}\otimes\overline{R}_{l,E}\right)\bigotimes_{i=1}^{N}|\phi^{+,d_{i}}_{A_{i}E_{i}}\rangle\!\langle\phi^{+,d_{i}}_{A_{i}E_{i}}|\right]$ (72)
where
$\overline{R}_{l,E}=D_{N}\left(\bigotimes_{i=1}^{N}P_{E_{i}}U_{E_{i}}\right)\ R_{l}\ \left(\bigotimes_{i=1}^{N}U_{E_{i}}^{\dagger}P_{E_{i}}\right),$ (73)
such that $D_{N}=\Pi_{i=1}^{N}d_{i}$. Now, notice that the state $\bigotimes_{i=1}^{N}|\phi^{+,d_{i}}_{A_{i}E_{i}}\rangle$ can also be expressed as a singe maximally entangled state between the external parties $A$ and $E$ of the local dimension $D_{N}$,
$\bigotimes_{i=1}^{N}|\phi^{+,d_{i}}_{A_{i}E_{i}}\rangle=|\phi^{+,D_{N}}_{AE}\rangle,$ (74)
where
$|\phi^{+,D_{N}}_{A|E}\rangle=\frac{1}{\sqrt{D_{N}}}\sum_{j=0}^{D_{N}-1}|j\rangle_{A}|j\rangle_{E}.$ (75)
Plugging this state back to the condition (72) and then using the well-known property of the maximally entangled state that $\mathbbm{1}_{A}\otimes Q_{B}|\phi^{+,D}_{AB}\rangle=Q^{T}_{A}\otimes\mathbbm{1}_{B}|\phi^{+,D}_{AB}\rangle$ for any matrix $Q$, we have that
$|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\tilde{\rho}^{l}_{A^{\prime\prime}}=2^{N}\mathrm{Tr}_{E}\left[\overline{R}^{T}_{l,A}\otimes\mathbbm{1}_{E}|\phi^{+,D_{N}}_{A|E}\rangle\!\langle\phi^{+,D_{N}}_{A|E}|\right].$ (76)
After taking the trace over $E$, the above formula simplifies to
$|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\tilde{\rho}^{l}_{A^{\prime\prime}}=\frac{2^{N}}{D_{N}}\overline{R}^{T}_{l,A}.$ (77)
We can now apply the transposition to both sides of the above relations and then expand its right-hand side using (73), to obtain
$|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\tilde{\rho}^{l,T}_{A^{\prime\prime}}=2^{N}\left(\bigotimes_{i=1}^{N}P_{A_{i}}U_{{A_{i}}}\right)\ R_{l}\ \left(\bigotimes_{i=1}^{N}U_{{A_{i}}}^{\dagger}P_{{A}_{i}}\right),$ (78)
which after taking the sum over $l$ and using the fact that the measurement operators $R_{l}$ sum up to the identity, leads further to
$\sum_{l=0}^{2^{N}-1}|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\tilde{\rho}^{l,T}_{A^{\prime\prime}}=2^{N}\bigotimes_{i=1}^{N}P_{A_{i}}^{2}.$ (79)
Now, taking a partial trace over the subsystems $A_{2},A_{3},\ldots,A_{N}$ in the above expression, we obtain
$\frac{\mathbbm{1}_{A_{1}^{\prime}}}{2}\otimes\sigma_{A_{1}^{\prime\prime}}=P_{A_{1}}^{2},$ (80)
which by virtue of the fact that $P_{A_{i}}\geq 0$ for any $i$, implies that
$\mathbbm{1}_{A_{1}^{\prime}}\otimes\sqrt{\frac{\sigma_{A_{1}^{\prime\prime}}}{2}}=P_{A_{1}}$ (81)
where,
$\sigma_{A_{1}^{\prime\prime}}=\frac{1}{2^{N}}\sum_{l=0}^{2^{N}-1}\mathrm{Tr}_{A_{2}^{\prime\prime}\ldots A_{N}^{\prime\prime}}\left(\tilde{\rho}^{l,T}_{A^{\prime\prime}}\right).$ (82)
In a similar way we can determine from Eq. (79) the form of the other matrices $P_{A_{i}}$ and thus
$P_{A_{i}}=\mathbbm{1}_{A_{i}^{\prime}}\otimes\sqrt{\frac{\sigma_{A_{i}^{\prime\prime}}}{2}}\qquad(i=1,\ldots,N),$ (83)
where,
$\sigma_{A_{j}^{\prime\prime}}=\frac{1}{2^{N}}\sum_{l=0}^{2^{N}-1}\mathrm{Tr}_{A^{\prime\prime}\setminus\{A_{j}^{\prime\prime}\}}\left(\tilde{\rho}^{l,T}_{A^{\prime\prime}}\right).$ (84)
Plugging $P_{A_{j}}$ in the state (70) and also recalling that $d_{i}=2d^{\prime}_{i}$, we obtain
$|\overline{\psi}_{A_{i}E_{i}}\rangle=\sqrt{d^{\prime}_{i}}\ \left(\mathbbm{1}_{A_{i}E_{i}^{\prime}}\otimes\sqrt{\sigma_{E_{i}^{\prime\prime}}}\right)|\phi^{+,d_{i}}_{A_{i}E_{i}}\rangle\qquad\forall i.$ (85)
Using then again the fact that $d_{i}$ is even, which implies that the maximally entangled state $|\phi^{+,d_{i}}_{A_{i}E_{i}}\rangle$ can also be written as
$|\phi^{+,d_{i}}_{A_{i}E_{i}}\rangle=|\phi^{+}_{A_{i}^{\prime}E_{i}^{\prime}}\rangle|\phi^{+,d^{\prime}_{i}}_{A_{i}^{\prime\prime}E_{i}^{\prime\prime}}\rangle,$ (86)
we finally conclude that there are unitary operations $U_{A_{i}}$ and $U_{E_{i}}$ such that
$(U_{A_{i}}\otimes U_{E_{i}})|\psi_{A_{i}E_{i}}\rangle=|\overline{\psi}_{A_{i}E_{i}}\rangle=|\phi^{+}_{A_{i}^{\prime}E_{i}^{\prime}}\rangle|\xi_{A_{i}^{\prime\prime}E_{i}^{\prime\prime}}\rangle\qquad\forall i.$ (87)
where the auxiliary state $|\xi_{A_{i}^{\prime\prime}E_{i}^{\prime\prime}}\rangle$ is given by,
$|\xi_{A_{i}^{\prime\prime}E_{i}^{\prime\prime}}\rangle=\left(\mathbbm{1}_{A_{i}^{\prime\prime}}\otimes\sqrt{d^{\prime}_{i}\sigma_{E_{i}^{\prime\prime}}}\right)|\phi^{+,d^{\prime}_{i}}_{A^{\prime\prime}_{i}E^{\prime\prime}_{i}}\rangle.$ (88)
C.  Entangled measurement $\{R_{l}\}$
Let us now concentrate on the measurement $\{R_{l}\}$. To this end, we exploit the explicit form of $P_{A_{i}}$ and fact that the matrices $\sigma_{A_{i}^{\prime\prime}}$ are invertible (which is a consequence of the assumption that the states produced by the sources are locally full rank) to rewrite Eq. (78) as
$\left(\bigotimes_{i=1}^{N}U_{{A_{i}}}\right)\ R_{l}\ \left(\bigotimes_{i=1}^{N}U_{{A_{i}}}^{\dagger}\right)=|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\left(\bigotimes_{i=1}^{N}\sigma^{-1/2}_{A_{i}^{\prime\prime}}\right)\tilde{\rho}^{l,T}_{A^{\prime\prime}}\left(\bigotimes_{i=1}^{N}\sigma^{-1/2}_{A_{i}^{\prime\prime}}\right),$ (89)
and further in a simpler way as
$\left(\bigotimes_{i=1}^{N}U_{{A_{i}}}\right)\ R_{l}\ \ \left(\bigotimes_{i=1}^{N}U_{{A_{i}}}^{\dagger}\right)=|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes(\tilde{R}_{l})_{A^{\prime\prime}},$ (90)
where the matrices $\tilde{R}_{l}$ acting on the $A_{i}^{\prime\prime}$ subsystems are defined as
$\tilde{R}_{l,A^{\prime\prime}}=\left(\bigotimes_{i=1}^{N}\sigma^{-1/2}_{A_{i}^{\prime\prime}}\right)\tilde{\rho}^{l,T}_{A^{\prime\prime}}\left(\bigotimes_{i=1}^{N}\sigma^{-1/2}_{A_{i}^{\prime\prime}}\right).$ (91)
Now, an important step is to observe that after implementing the fact that $\sum_{l}R_{l}=\mathbbm{1}$, Eq. (90) allows us to conclude that
$\mathbbm{1}_{A}=\sum_{l=0}^{2^{N}-1}|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes(\tilde{R}_{l})_{A^{\prime\prime}}.$ (92)
Then, since the vectors $|\phi_{l}\rangle$ are mutually orthogonal, the above directly implies that
$\tilde{R}_{l}=\mathbbm{1}_{A^{\prime\prime}}$ (93)
using which we finally obtain from Eq. (90) that the measurement operators $R_{l}$ are up to a local unitary transformation given by
$\left(\bigotimes_{i=1}^{N}U_{{A_{i}}}\right)\ R_{l}\ \ \left(\bigotimes_{i=1}^{N}U_{{A_{i}}}^{\dagger}\right)=|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\mathbbm{1}_{A^{\prime\prime}}\qquad\forall l.$ (94)
D.  Measurements $A_{i,2}$
Let us finally provide a self-testing statement for the third measurement of the external parties $A_{i,2}$ for any $i$. For this purpose, we consider Eq. (28c) and substitute into it the certified observables and states from Eqs. (3) and (65) to obtain a set of equations
$\left[-(-1)^{l_{i}}\overline{A}_{1,2}\overline{A}_{i,2}\otimes\bigotimes_{\begin{subarray}{c}j=2\\ j\neq i\end{subarray}}^{N}X_{A_{j}^{\prime}}\right]|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\tilde{\rho}^{l}_{A^{\prime\prime}}=|\phi_{l}\rangle\!\langle\phi_{l}|_{A^{\prime}}\otimes\tilde{\rho}^{l}_{A^{\prime\prime}}\qquad i=2,\ldots,N.$ (95)
where $\overline{A}_{i,2}=U_{A_{i}}A_{i,2}\,U_{A_{i}}^{\dagger}$. As the local Hilbert spaces are even-dimensional, we can express the observables $\overline{A}_{i,2}$ as
$\overline{A}_{i,2}=\mathbbm{1}\otimes P_{i}+Z\otimes Q_{i}+X\otimes R_{i}+Y\otimes S_{i},$ (96)
where $P_{i},Q_{i},R_{i},S_{i}$ are hermitian matrices that mutually commute and satisfy $P_{i}^{2}+Q_{i}^{2}+R_{i}^{2}+S_{i}^{2}=\mathbbm{1}$ for any $i$.
Let us first look at the case when $N=2$. Putting the form of the measurements (96) into the relation (95) and then sandwiching it with $|\phi_{l}\rangle$, we obtain
$\left[P_{1}\otimes P_{2}+(-1)^{l_{1}+l_{2}}Q_{1}\otimes Q_{2}+(-1)^{l_{1}}R_{1}\otimes R_{2}-(-1)^{l_{2}}S_{1}\otimes S_{2}\right]\tilde{\rho}^{l}_{A_{1}^{\prime\prime}A_{2}^{\prime\prime}}=-(-1)^{l_{2}}\tilde{\rho}^{l}_{A_{1}^{\prime\prime}A_{2}^{\prime\prime}}$ (97)
for any $l_{1},l_{2}=0,1$. Let us now observe that from Eq. (91) and the fact that $\tilde{R}_{l}=\mathbbm{1}_{A^{\prime\prime}}$ it follows that the matrices $\tilde{\rho}^{l}_{A_{1}^{\prime\prime}A_{2}^{\prime\prime}}$ are full rank and therefore the above implies the following four matrix equations
$P_{1}\otimes P_{2}+(-1)^{l_{1}+l_{2}}Q_{1}\otimes Q_{2}+(-1)^{l_{1}}R_{1}\otimes R_{2}-(-1)^{l_{2}}S_{1}\otimes S_{2}=-(-1)^{l_{2}}\mathbbm{1},$ (98)
we can directly be solved to obtain
$S_{1}\otimes S_{2}=\mathbbm{1},\qquad P_{1}\otimes P_{2}=Q_{1}\otimes Q_{2}=R_{1}\otimes R_{2}=0.$ (99)
Now, considering the general case $N\geq 3$ and again putting form of the measurements (98) into (95), and sandwiching the above expression with the state $|\phi_{l}\rangle$, we obtain for any $l_{i}=0,1$
$\left[(-1)^{l_{1}}R_{1}\otimes R_{i}-(-1)^{l_{i}}S_{1}\otimes S_{i}\right]\tilde{\rho}^{l}_{A^{\prime\prime}}=-(-1)^{l_{i}}\tilde{\rho}^{l}_{A^{\prime\prime}}\qquad(i=2,\ldots,N)$ (100)
As $\tilde{\rho}^{l}_{A^{\prime\prime}}$ is full-rank, the above translates to
$(-1)^{l_{1}}R_{1}\otimes R_{i}-(-1)^{l_{i}}S_{1}\otimes S_{i}=-(-1)^{l_{i}}\mathbbm{1}.$ (101)
For any $i$, we get two different equations corresponding to two values of $l_{1}=0$ and $l_{1}=1$, which directly imply
$S_{1}\otimes S_{i}=\mathbbm{1},\qquad R_{1}\otimes R_{i}=0\qquad(i=2,\ldots,N).$ (102)
Now, one notices that $S_{1}^{2}\otimes S_{i}^{2}=\mathbbm{1}$ which together with the upper bound $S_{i}^{2}\leq\mathbbm{1}$ which follows from the condition $P_{i}^{2}+Q_{i}^{2}+R_{i}^{2}+S_{i}^{2}=\mathbbm{1}$, allows us to conclude that $S_{i}^{2}=\mathbbm{1}$ for any $i$. This also means that $P^{2}_{i}=R^{2}_{i}=Q^{2}_{i}=0$ and consequently that
$P_{i}=R_{i}=Q_{i}=0.$ (103)
This analysis leads us to the conclusion that $\tilde{A}_{i,2}=Y\otimes S_{i}$ with $S_{i}^{2}=\mathbbm{1}$. It turns out, however, that the conditions (102) are more restrictive. To see that explicitly, let us follow the reasoning applied already in Ref. [20] and observe that each $S_{i}$ can be decomposed us $S_{i}=S_{i}^{+}-S_{i}^{-}$, where $S_{i}^{\pm}$ are projections onto the eigensubspaces of $S_{i}$ corresponding to the eigenvalues ±$\pm 1$ with $S_{i}^{+}+S_{i}^{-}=\mathbbm{1}$. After plugging this representation into $S_{1}\otimes S_{i}=\mathbbm{1}$ and rearranging terms one arrives at the following conditions
$S_{1}^{+}\otimes S_{i}^{-}=S_{1}^{-}\otimes S_{i}^{+}=0\qquad(i=2,\ldots,N).$ (104)
These conditions imply that there are two possible solutions for Alice’s third measurements: either
$\tilde{A}_{i,2}=Y_{A_{i}^{\prime}}\otimes\mathbbm{1}_{A_{i}^{\prime\prime}}\qquad(i=1,\ldots,N)$ (105)
or
$\tilde{A}_{i,2}=-Y_{A_{i}^{\prime}}\otimes\mathbbm{1}_{A_{i}^{\prime\prime}}\qquad(i=1,\ldots,N).$ (106)
This completes the proof. ∎
Appendix B Appendix B: Self-testing any projective measurement
In this section, we show how to certify Eve’s second measurement $E_{1}=\{R_{l|1}\}$ to be a particular reference projective measurement $E^{\prime}_{1}=\{R^{\prime}_{l|1}\}$ (up to some additional degrees of freedom). For this purpose, we use the states generated by the sources, certified as in Eq. (30), and the certified measurements in Eqs. (31) and (32) along with some additional conditions imposed on the observed correlations.
Let us consider an ideal reference projective measurement $E^{\prime}_{1}=\{R^{\prime}_{l|1}\}$ where $R^{\prime}_{l|1}$ are mutually orthogonal projections defined on a Hilbert space of arbitrary finite dimension $\mathbbm{C}^{D}$ whose ranks are in general arbitrary; let denote them by $r_{l}=\mathrm{rank}(R^{\prime}_{l|1})$. Let us also assume that the measurement has $K\leq 2^{N}$ outcomes; $r_{1}+\ldots+r_{K}=D$. Such a measurement can always be embedded in an $N$-qubit Hilbert space $(\mathbbm{C}^{2})^{\otimes N}$ where $N$ the minimal natural number such that $D\leq 2^{N}$. In order to represent $R^{\prime}_{l|1}$ in the $N$-qubit Hilbert space one can for instance assign to any element of the standard basis $|i\rangle\in\mathbbm{C}^{D}$ an element $|i_{1}\ldots i_{N}\rangle$ of the product standard basis of $(\mathbbm{C}^{2})^{\otimes}$ such that $i_{1}\ldots i_{N}$ is a binary representation of $i$ $(i=0,\ldots,D-1)$. Then, one needs to add another projective measurement element $M^{\perp}$ to $E_{1}^{\prime}$ so that the resulting measurement defined on $(\mathbbm{C}^{2})^{\otimes N}$ satisfies
$\sum_{i=1}^{K}R^{\prime}_{i|1}+M^{\perp}=\mathbbm{1}_{2^{N}},$ (107)
$\mathbbm{1}_{2^{N}}$ is the identity acting on $(\mathbbm{C}^{2})^{\otimes N}$. Given that the above procedure applies to any projective measurement, We can assume that the reference projective measurement $E_{1}^{\prime}$ is from the beginning defined on the $N$-qubit Hilbert space.
Let us then decompose every measurement element in the Pauli basis,
$\sigma_{0}=Z,\quad\sigma_{1}=X,\quad\sigma_{2}=Y,\quad\sigma_{3}=\mathbbm{1}_{2},$ (108)
as
$R^{\prime}_{l|1}=\sum_{i_{1},\ldots,i_{N}=0}^{3}f^{l}_{i_{1},\ldots,i_{N}}\sigma_{1,i_{1}}\otimes\ldots\otimes\sigma_{N,i_{N}},$ (109)
where the additional subscript marks the system on which the Pauli matrix $\sigma_{k,i_{k}}$ acts; for instance, $\sigma_{3,2}$ is $\sigma_{2}$ acting on site $3$. Then, $f_{i_{1},\ldots,i_{N}}^{l}$ are real numbers defined as
$f_{i_{1},\ldots,i_{N}}^{l}=\frac{1}{2^{N}}\mathrm{Tr}[(\sigma_{1,i_{1}}\otimes\ldots\otimes\sigma_{N,i_{N}})R^{\prime}_{l|1}].$ (110)
Let us now move on to our self-testing scheme and consider again the quantum network scenario presented in above. Let us then assume that apart from the conditions imposed in Theorem 1, the correlations observed in the network satisfy the following additional set of constraints
$\sum_{i_{1},\ldots,i_{N}=0}^{3}f^{l}_{i_{1},\ldots,i_{N}}\left\langle\tilde{A}_{1,i_{1}}\otimes\bigotimes_{k=2}^{N}A_{k,i_{k}}\otimes(R_{l|1})_{E}\right\rangle_{|\psi_{AE}\rangle}$ $=$ $\frac{r_{l}}{2^{N}}\quad\forall l,$ (111)
where $\tilde{A}_{1,0},\tilde{A}_{1,1}$ are given in (18) with $\tilde{A}_{1,2}=A_{1,2},\tilde{A}_{1,3}=\mathbbm{1}$ and $A_{k,3}=\mathbbm{1}$ for any $k$.
We can now present our result.
Theorem 2.
Let us suppose that the correlations generated in the quantum network satisfy the assumptions of Theorem 1 and the additional conditions in Eq. (111). Then, for any $l=1,\ldots,K$ we conclude that
[…]
$U_{E}R_{l|1}\,U_{E}^{\dagger}$ $=$ $R^{\prime}_{l|1}\otimes\mathbbm{1}_{E^{\prime\prime}}\quad\ \ \ \mathrm{if}\quad A_{i,2}=-Y\otimes\mathbbm{1}\quad\forall i$ (112)
where $U_{E}$ is the same unitary as in Theorem 1 and $E^{\prime\prime}=E_{1}^{\prime\prime}\ldots E_{N}^{\prime\prime}$.
Proof.
For simplicity, we represent $R_{l|1}\equiv R_{l}$ throughout the proof. Let us first consider the relation in Eq. (111) for a particular $l$ and then expand it by using the fact that the observables $A_{s,j}$ for $s=1,\ldots,N$ and $j=0,1,2$ are certified as in Eqs. (31) and (32). As $A_{s,2}$ for any $s$ can have two realisations as given in Eq. (32), we first consider the case when $A_{s,2}=Y\otimes\mathbbm{1}_{s^{\prime\prime}}$. This gives us
$\sum_{i_{1},\ldots,i_{N}=0}^{3}f^{l}_{i_{1},\ldots,i_{N}}\left\langle\left(\bigotimes_{s=1}^{N}U_{A_{s}}^{\dagger}\right)\left(\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}\otimes\mathbbm{1}_{A^{\prime\prime}}\right)\left(\bigotimes_{s=1}^{N}U_{A_{s}}\right)\otimes R_{l}\right\rangle_{\psi_{AE}}=\frac{r_{l}}{2^{N}}.$ (113)
Then, it follows from Theorem 1 [cf. Eq. (29)] that the global state can be represented in the following way
$|\psi_{AE}\rangle=\bigotimes_{s=1}^{N}|\psi_{A_{s}E_{s}}\rangle=\bigotimes_{s=1}^{N}(U_{A_{s}}^{\dagger}\otimes U_{E_{s}}^{\dagger})(|\phi^{+}_{A_{s}^{\prime}E_{s}^{\prime}}\rangle\otimes|\xi_{A_{s}^{\prime\prime}E_{s}^{\prime\prime}}\rangle).$ (114)
Notice also that by virtue of Eq. (88) the junk states $|\xi_{A_{s}^{\prime\prime}E_{s}^{\prime\prime}}\rangle$ can be represented as
$|\xi_{A_{i}^{\prime\prime}E_{i}^{\prime\prime}}\rangle=\left(\mathbbm{1}_{A_{i}^{\prime\prime}}\otimes\sqrt{d^{\prime}_{i}\sigma_{E_{i}^{\prime\prime}}}\right)|\phi^{+}_{A^{\prime\prime}_{i}E^{\prime\prime}_{i}}\rangle,$ (115)
where we omitted the local dimension $d_{i}^{\prime}$ of the maximally entangled state. Thus, the joint state can further be represented in terms of a single maximally entangled states between $A$ and $E$ systems [cf. Eq. (74)] as
$\left(\bigotimes_{s=1}^{N}U_{A_{s}}\otimes U_{E_{s}}\right)|\psi_{AE}\rangle=\left(\bigotimes_{s=1}^{N}\sqrt{d^{\prime}_{s}\sigma_{E_{s}^{\prime\prime}}}\right)|\phi^{+}_{D_{N}}\rangle_{A|E}.$ (116)
Here the local dimension of the state is $D_{N}=d_{1}\ldots d_{N}=2^{N}d_{1}^{\prime}\ldots d_{N}^{\prime}$ and we used $A|E$ to highlight the fact that $|\phi^{+}_{D_{N}}\rangle_{A|E}$ is a maximally entangled state between all external parties $A_{i}$ and the central one $E$.
Plugging the state (116) into Eq. (113), and then using the fact that $(\mathbbm{1}\otimes Q)|\phi^{+}_{d}\rangle=(Q^{T}\otimes\mathbbm{1})|\phi^{+}_{d}\rangle$ for any matrix $Q$, we obtain
$\sum_{i_{1},\ldots,i_{N}=0}^{3}f^{l}_{i_{1},\ldots,i_{N}}\left\langle\left(\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}\otimes\bigotimes_{s=1}^{N}d^{\prime}_{s}\sigma^{T}_{A_{s}^{\prime\prime}}\right)\overline{R}_{l}^{T}\otimes\mathbbm{1}_{E}\right\rangle_{\phi^{+}_{D_{N}}}=\frac{r_{l}}{2^{N}},$ (117)
where
$\overline{R}_{l}=\left(\bigotimes_{s=1}^{N}U_{E_{s}}\right)R_{l}\left(\bigotimes_{s=1}^{N}U_{E_{s}}^{\dagger}\right).$ (118)
Then, tracing out the Eve’s subsystems from the expectation value in Eq. (117), we arrive at
$\sum_{i_{1},\ldots,i_{N}=0}^{3}f^{l}_{i_{1},\ldots,i_{N}}\mathrm{Tr}\left[\left(\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}\otimes\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}\right)\overline{R}_{l}^{T}\right]=r_{l}.$ (119)
Using Eq. (109), we can simplify the above formula as
$\mathrm{Tr}\left[\left(R^{\prime}_{l|1}\otimes\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}\right)\overline{R}_{l}^{T}\right]=r_{l}.$ (120)
Let us now concentrate on a particular outcome $l$ and consider an orthonormal basis in $\mathbbm{C}^{2^{N}}$ which contains the eigenvectors of the projection $R^{\prime}_{l|1}$; we denote this basis by =$\mathcal{B}=\{|\delta_{m}\rangle\}$, where the first $r_{l}$ vectors for $m=1,\ldots,r_{l}$ are the eigenvectors of $R^{\prime}_{l|1}$. Due to the fact that $\overline{R}_{0}^{T}$ acts on the Hilbert space $(\mathbbm{C}^{2})^{\otimes N}\otimes\bigotimes_{s=1}^{N}\mathcal{H}_{{A_{s}^{\prime\prime}}}$, we can express it using the basis $\mathcal{B}$ as
$\overline{R}_{0}^{T}=\sum_{m,m^{\prime}}|\delta_{m}\rangle\!\langle\delta_{m^{\prime}}|\otimes\tilde{R}_{m,m^{\prime}},$ (121)
where $\tilde{R}_{l,l^{\prime}}$ are some unknown matrices acting on $\bigotimes_{s=1}^{N}\mathcal{H}_{{A_{s}^{\prime\prime}}}$.
After plugging Eq. (121) into (120) and using the cyclic property of trace, we then obtain
$\sum_{m=1}^{r_{l}}\mathrm{Tr}\left[\left(\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}\right)\tilde{R}_{m,m}\right]=r_{l}.$ (122)
Since each term under the above sum is upper-bounded by one, we have that
$\mathrm{Tr}\left[\left(\bigotimes_{s=1}^{N}\sigma_{A_{s}^{\prime\prime}}^{T}\right)\tilde{R}_{m,m}\right]=1\qquad(m=1,\ldots,r_{l}).$ (123)
Now, due to the fact that the density matrices $\sigma_{A^{\prime\prime}_{s}}$ are full rank and $0\leqslant\tilde{R}_{m,m}\leqslant\mathbbm{1}$ which stems from the fact that $\{R_{l}\}$ is a quantum measurement, one observers that the above equation holds true if, and only if, $\tilde{R}_{m,m}=\mathbbm{1}$ for any $m=1,\ldots,r_{0}$. This implies that
$\overline{R}_{l}^{T}$ $=$ $R^{\prime}_{l|1}\otimes\mathbbm{1}_{A^{\prime\prime}}+\mathbbm{L}_{l},$ (124)
where $\mathbbm{L}_{l}$ is an operator composed of the remaining terms of $\overline{R}_{l}$ appearing in the decomposition (121); in other words, $\mathbbm{L}_{l}=\overline{R}_{l}^{T}-R^{\prime}_{l|1}\otimes\mathbbm{1}$.
$\mathbbm{L}_{l}$ $=$ $\sum_{\begin{subarray}{c}m,m^{\prime}=1\\ m\neq m^{\prime}\end{subarray}}^{r_{l}}|\delta_{m}\rangle\!\!\langle\delta_{m^{\prime}}|\otimes\tilde{R}_{m,m^{\prime}}+\sum_{m=1}^{r_{l}}\sum_{m^{\prime}=r_{l}+1}^{2^{N}}|\delta_{m}\rangle\!\!\langle\delta_{m^{\prime}}|\otimes\tilde{R}_{m,m^{\prime}}$ (128)
$+\sum_{m=r_{l}+1}^{2^{N}}\sum_{m^{\prime}=1}^{r_{l}}|\delta_{m}\rangle\!\!\langle\delta_{m^{\prime}}|\otimes\tilde{R}_{m,m^{\prime}}+\sum_{m=r_{l}+1}^{2^{N}}\sum_{m^{\prime}=r_{l}+1}^{2^{N}}|\delta_{m}\rangle\!\!\langle\delta_{m^{\prime}}|\otimes\tilde{R}_{m,m^{\prime}}.$
Let us now show that the fact that $\overline{R}_{l}^{T}\leq\mathbbm{1}$ imposes that the first three sums in the above representation must vanish. To this aim, let us consider a product vector $|\delta_{m}\rangle|\xi\rangle$ for any $m=1,\ldots,r_{l}$ and any vector $|\xi\rangle$, and act on it with $\overline{R}_{l}^{T}=R^{\prime}_{l|1}\otimes\mathbbm{1}_{A^{\prime\prime}}+\mathbbm{L}_{l}$. This gives
$\overline{R}_{l}^{T}|\delta_{m}\rangle|\xi\rangle=|\delta_{m}\rangle|\xi\rangle+\sum_{\begin{subarray}{c}m^{\prime}=1\\ m^{\prime}\neq m\end{subarray}}^{2^{N}}|\delta_{m^{\prime}}\rangle\otimes\tilde{R}_{m^{\prime},m}|\xi\rangle,$ (129)
which further leads to the following condition
$\langle\delta_{m}|\langle\xi|(\overline{R}_{l}^{T})^{2}|\delta_{m}\rangle|\xi\rangle=1+\sum_{\begin{subarray}{c}m^{\prime}=1\\ m^{\prime}\neq m\end{subarray}}^{2^{N}}\langle\xi|\tilde{R}_{m^{\prime},m}^{\dagger}\tilde{R}_{m^{\prime},m}|\xi\rangle.$ (130)
Let us observe now that $\overline{R}_{l}^{T}\leq\mathbbm{1}$ implies $(\overline{R}_{l}^{T})^{2}\leq\mathbbm{1}$, which after applying to the above equation allows one to conclude that
$\sum_{\begin{subarray}{c}m^{\prime}=1\\ m^{\prime}\neq m\end{subarray}}^{2^{N}}\langle\xi|\tilde{R}_{m^{\prime},m}^{\dagger}\tilde{R}_{m^{\prime},m}|\xi\rangle\leq 0.$ (131)
Therefore, $\langle\xi|\tilde{R}_{m^{\prime},m}^{\dagger}\tilde{R}_{m^{\prime},m}|\xi\rangle=0$ for any vector $|\xi\rangle$, which in turn means that $\tilde{R}_{m^{\prime},m}=0$ for any $m^{\prime}\neq m$. Since the same argument can be repeated for any $|\delta_{m}\rangle$ with $m=1,\ldots,r_{l}$ one then has that all the matrices $\tilde{R}_{m^{\prime},m}=0$ for any $m=1,\ldots,r_{l}$ and any $m^{\prime}\neq m$. This finally implies that the $\mathbbm{L}_{l}$ operator can be rewritten as
$\mathbbm{L}_{l}=\sum_{m=r_{l}+1}^{2^{N}}\sum_{m^{\prime}=r_{l}+1}^{2^{N}}|\delta_{m}\rangle\!\langle\delta_{m^{\prime}}|\otimes\tilde{R}_{m,m^{\prime}},$ (132)
and thus it acts on a subspace of the corresponding Hilbert space which is orthogonal to that supporting $R^{\prime}_{l|1}\otimes\mathbbm{1}_{A^{\prime\prime}}$. Due to the fact that the measurement operator satisfies $\overline{R}_{l}\geq 0$ the latter clearly implies that $\mathbbm{L}_{l}\geq 0$.
It is important to stress that the same reasoning can be applied to $\overline{R}_{l}^{T}$ for any $l$, and thus every $\overline{R}_{l}^{T}$ decomposes into a direct sum
$\overline{R}_{l}^{T}=R^{\prime}_{l|1}\otimes\mathbbm{1}_{A^{\prime\prime}}+\mathbbm{L}_{l},$ (133)
of two positive semi-definite operators $R^{\prime}_{l|1}\otimes\mathbbm{1}_{A^{\prime\prime}}$ and $\mathbbm{L}_{l}$. Now, after summing (133) over all outcomes, one obtains
$\sum_{l}\mathbbm{L}_{l}=0,$ (134)
where we also exploited the fact that both $\overline{R}_{l}$ and $R^{\prime}_{l|1}$ sum up to the identities. Taking into account that $\mathbbm{L}_{l}\geq 0$, one deduces from the condition (134) that $_{l}=0$ for any $l$ and thus
$\overline{R}_{l}^{T}=R^{\prime}_{l|1}\otimes\mathbbm{1}_{A^{\prime\prime}}$ (135)
or, equivalently,
$\overline{R}_{l}=(R^{\prime}_{l|1})^{*}\otimes\mathbbm{1}_{A^{\prime\prime}}.$ (136)
Similarly, one can consider the other case of $A_{s,2}=-Y\otimes\mathbbm{1}_{s^{\prime\prime}}$ for any $s$ and obtain an analogous relation to Eq. (120) which reads
$\mathrm{Tr}\left[\left((R^{\prime}_{l|1})^{*}\otimes\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}\right)\overline{R}_{l}^{T}\right]=r_{l},\qquad(l=1,\ldots,K).$ (137)
Following exactly the same steps as above, we find that $\overline{R}_{l}=R^{\prime}_{l|1}\otimes\mathbbm{1}$ for $l=1,\ldots,K$, thus completing the proof. ∎
Appendix C Appendix C: Self-testing any rank-one extremal quantum measurement
In this section, we show that one can self-test the measurement corresponding to the input $e=1$ with the central party to be a general extremal rank-one quantum measurement $E_{1}=\{R^{\prime}_{l|1}\}$ up to some additional degrees of freedom. For this purpose, we use the states generated by the sources $P_{i}$, certified as in Eq. (29), and the certified measurements in Eqs. (31) and (32) along with some additional conditions.
Let us then consider an ideal reference measurement $E^{\prime}_{1}=\{R^{\prime}_{l|1}\}$ with $R^{\prime}_{l|1}$ $(l=1,\ldots,K)$ being measurement elements defined on some Hilbert space $\mathbbm{C}^{D}$ and assume this measurement to be extremal. For completeness let us recall that an extremal POVM is one that cannot be decomposed into a convex combination of other POVM’s. Additionally, we assume that the measurement elements are rank-one, in which case they can be expressed as
$R^{\prime}_{l|1}=\lambda_{l}|\delta_{l}\rangle\!\langle\delta_{l}|,$ (138)
where $0\leqslant\lambda_{l}\leqslant 1$ for any $l$ with the property that the states $\{|\delta_{l}\rangle\}$ are linearly independent [cf. Ref. [34]].
Let us now notice that similarly to the case of projective measurements discussed above, the extremal rank-one POVM $\{R^{\prime}_{l|1}\}$ can also be embedded in $(\mathbbm{C}^{2})^{\otimes N}$. To this end we consider the minimal $N$ such that $D\leq 2^{N}$ and associate to every element of the standard basis $|i\rangle\in\mathbbm{C}^{D}$ a vector $|i_{1}\ldots i_{N}\rangle$ forming the standard basis in $\in(\mathbbm{C}^{2})^{\otimes N}$ such that $i_{1}\ldots i_{N}$ is the binary representation of $i=0,\ldots,D-1$. Since then the measurement elements do not sum up to the identity on $(\mathbbm{C}^{2})^{\otimes N}$, we need to consider additional measurement elements $J_{k}$ $(k=1,\ldots,2^{N}-D)$ such that
$\sum_{l=1}^{K}R^{\prime}_{l|1}+\sum_{k=1}^{2^{N}-D}J_{k}=\mathbbm{1}_{2^{N}}$ (139)
We take this additional operators to be rank-one projections, $J_{k}=|\varphi_{k}\rangle\!\langle\varphi_{k}|$ where $|\varphi_{k}\rangle$ are mutually orthogonal. Moreover, the vectors $|\varphi\rangle$ are orthogonal to $\sum_{l=1}^{K}R^{\prime}_{l|1}$. Such vectors can always be found, for instance by taking an arbitrary orthonormal basis in the subspace of $(\mathbbm{C}^{2})^{\otimes}$ which is orthogonal to the support of $\sum_{l=1}^{K}R^{\prime}_{l|1}$. Let us denote the new measurement by $N=\{\lambda_{1}|\delta_{1}\rangle\!\langle\delta_{1}|,\ldots,\lambda_{K}|\delta_{K}\rangle\!\langle\delta_{K}|,J_{1},\ldots,J_{2^{N}-D}\}$ and notice that it is extremal too because it is obtained by complementing the extremal measurement $E_{1}^{\prime}$ by rank-one projections [34].
Let us now represent any measurement element $N_{l}$ in terms of the Pauli basis (108) as
$N_{l}=\sum_{i_{1},\ldots,i_{N}=0}^{3}f^{l}_{i_{1},\ldots,i_{N}}\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}.$ (140)
Let us again consider the quantum network scenario presented and assume that the observed correlations additionally satisfy the following conditions
$\left\langle\tilde{A}_{1,i_{1}}\otimes\bigotimes_{k=2}^{N}A_{k,i_{k}}\otimes(R_{l|1})_{E}\right\rangle_{\psi_{AE}}$ $=$ $f^{l}_{i_{1},\ldots,i_{N}}\quad\forall l,i_{1},\ldots,i_{N},$ (141)
where $\tilde{A}_{1,0}$ and $\tilde{A}_{1,1}$ are given in (18) with $\tilde{A}_{1,2}=A_{1,2}$, $\tilde{A}_{1,3}=\mathbbm{1}$ and $A_{k,3}=\mathbbm{1}$ for any $k$.
Notice that the above statistics can be realised if the sources generate the two-qubit maximally entangled state $|\phi^{+}\rangle=(1/\sqrt{2})(|00\rangle+|11\rangle)$ and the measurement $E_{1}=\{R_{l|1}\}$ is exactly $R_{l|1}=(R^{\prime}_{l|1})^{*}$ where $R^{\prime}_{l|1}$ is given in (138) and *** denotes the complex conjugate, whereas the external parties perform the measurements in Eq. (19).
Next, we show that the above conditions along with certification of the states and measurements presented in Theorem 1 are sufficient to show that up to the standard equivalences and complex conjugation the unknown measurement $E_{1}=\{R_{l|1}\}$ is equivalent to the ideal one $E_{1}^{\prime}\equiv N$.
Theorem 3.
Let us suppose that the correlations generated in the quantum network satisfy the assumptions of Theorem 1 along with the additional constraints in Eq. (141). Then, for any $l=1,\ldots,K$ we can conclude that
[…]
$U_{E}R_{l|1}U_{E}^{\dagger}$ $=$ $R^{\prime}_{l|1}\otimes\mathbbm{1}_{E^{\prime\prime}}\quad\ \ \ \mathrm{if}\quad A_{i,2}=-Y\otimes\mathbbm{1}\quad\forall i$ (142)
where $U_{E}$ is the same unitary as in Theorem 1 and $E^{\prime\prime}=E_{1}^{\prime\prime}\ldots E_{N}^{\prime\prime}$.
Proof.
For simplicity, we represent $R_{l|1}\equiv R_{l}$ throughout the proof. Let us consider the relation in Eq. (141) and then substitute into it the observables $A_{s,j}$ for $s=1,\ldots,N$ and $j=0,1,2$ given in Eqs. (31) and (32). As $A_{s,2}$ for any $s$ can have two realisations as given in Eq. (32), we first consider the case when $A_{s,2}=Y\otimes\mathbbm{1}_{s^{\prime\prime}}$. This gives us
$\left\langle\left(\bigotimes_{s=1}^{N}U_{A_{s}}^{\dagger}\right)\left(\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}\otimes\mathbbm{1}_{A^{\prime\prime}}\right)\left(\bigotimes_{s=1}^{N}U_{A_{s}}\right)\otimes R_{l}\right\rangle_{\psi_{AE}}=f^{l}_{i_{1},\ldots,i_{N}}.$ (143)
It then follows from Theorem 1 that the global state $|\psi_{AE}\rangle$ can be represented as in Eq. (116), which allows us to rewrite the above condition in the following form
$d_{1}^{\prime}\ldots d_{N}^{\prime}\left\langle\left(\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}\otimes\bigotimes_{s=1}^{N}\sigma_{A_{s}^{\prime\prime}}^{T}\right)\overline{R}_{l}^{T}\otimes\mathbbm{1}_{E}\right\rangle_{\phi^{+}_{2^{N}D}}=f^{l}_{i_{1},\ldots,i_{N}},$ (144)
where we have also used the fact that $(\mathbbm{1}\otimes Q)|\phi^{+}_{d}\rangle=(Q^{T}\otimes\mathbbm{1})|\phi^{+}_{d}\rangle$ and $\overline{R}_{l}$ are given by
$\overline{R}_{l}=\left(\bigotimes_{s=1}^{N}U_{\overline{A}_{s}}\right)R_{l}\left(\bigotimes_{s=1}^{N}U_{\overline{A}_{s}}^{\dagger}\right).$ (145)
After tracing Eve’s subsystems the expression in Eq. (144) can further be simplified to
$\mathrm{Tr}\left[\left(\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}\otimes\bigotimes_{s=1}^{N}\sigma_{A_{s}^{\prime\prime}}^{T}\right)\overline{R}_{l}^{T}\right]=2^{N}f^{l}_{i_{1},\ldots,i_{N}}.$ (146)
Now, we can exploit the fact that $\overline{R}_{l}^{T}$ acts on the Hilbert space $\mathbbm{C}^{2^{N}}\otimes\bigotimes_{s=1}^{N}\mathcal{H}_{{A_{s}^{\prime\prime}}}$, we can again express it using the Pauli basis as
$\overline{R}_{l}^{T}=\sum_{i_{1},\ldots,i_{N}=0}^{3}\bigotimes_{k=1}^{N}\sigma_{k,i_{k}}\otimes\tilde{R}^{l}_{i_{1},\ldots,i_{N}}$ (147)
where $\tilde{R}_{i_{1},\ldots,i_{N}}$ acts on $\bigotimes_{s=1}^{N}\mathcal{H}_{{A_{s}^{\prime\prime}}}$. Then, using the fact that Pauli matrices are orthogonal in the Hilbert-Schmidt scalar product, $\mathrm{Tr}[\sigma_{i}\sigma_{j}]=2\delta_{ij}$ with $\delta_{ij}$ being the Kronecker’s symbol, we obtain
$\mathrm{Tr}\left[\left(\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}\right)\tilde{R}^{l}_{i_{1},\ldots,i_{N}}\right]=f^{l}_{i_{1},\ldots,i_{N}}\qquad\forall i_{1},\ldots,i_{N}.$ (148)
The rest of the proof follows exactly from Eqs. E13-E29 of [35]. One first needs to decompose the state $\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}$ in its eigenbasis, let’s say $\{|k\rangle\}$, and then use the fact that extremal POVM’s can not be decomposed into other POVM’s to obtain that
$\langle k|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|k\rangle=f^{l}_{i_{1},\ldots,i_{N}}\qquad\forall k,i_{1},\ldots,i_{N}.$ (149)
Then, using the fact that the measurement elements of extremal POVM’s are linearly independent, one can conclude that for all $i_{1},\ldots,i_{N}$
$\langle s|\tilde{R}^{l}_{i_{1},\ldots,i_{N}}|k\rangle=0\qquad\forall s,k\quad\mathrm{s.t.}\quad s\neq k$ (150)
which allows us to conclude that over the support of the state $\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}$
$\tilde{R}^{l}_{i_{1},\ldots,i_{N}}=f^{l}_{i_{1},\ldots,i_{N}}\mathbbm{1}_{E^{\prime\prime}}$ (151)
Similarly, when $A_{i,2}=-Y\otimes\mathbbm{1}$ for all $i$, then we will obtain exactly the similar relation to Eq. (148) as
$\mathrm{Tr}\left[\left(\bigotimes_{s=1}^{N}\sigma^{T}_{A_{s}^{\prime\prime}}\right)\tilde{R}_{l,i_{1},\ldots,i_{N}}\right]=g(i_{1},\ldots,i_{N})f_{l,i_{1},\ldots,i_{N}}\quad\forall i_{1},\ldots,i_{N}$ (152)
where $g(i_{1},\ldots,i_{N})=1$ for even number of $j$ and $g(i_{1},\ldots,i_{N})=-1$ for odd number of $j$ such that $i_{j}=2$. Thus, we can certify any extremal POVM. ∎
Appendix D Appendix D: Self-testing any quantum state
Using the Theorems 2 and 3 proven above, we can now show how it is possible to probabilistically self-test arbitrary pure or mixed $N$-qubit quantum state. For this purpose, let us first certify the post-measurement state when eve chooses $e=1$ and obtains an outcome $l$.
Corollary 1.
Assume that the states are certified as in Eq. (30) and Eve’s measurement $E_{1}$ is certified as in Eqs. (2) and (3). Consequently, when Eve observes the $l-$th outcome of her measurement, the post-measurement state with the external parties is given by
[…]
$U_{A}\,\rho_{A}^{l}\,U_{A}^{\dagger}$ $=$ $\frac{1}{\mathrm{Tr}R^{\prime}_{l|1}}(R^{\prime}_{l|1})^{*}\otimes\tilde{\rho}_{A^{\prime\prime}}\quad\ \ \ \mathrm{if}\quad A_{i,2}=-Y\otimes\mathbbm{1}\quad\forall i,$ (153)
where $U_{A}=\bigotimes_{s=1}^{N}U_{A_{s}}$ and the unitaries $U_{A_{s}}$ are the same as in Eq. (30).
Proof.
The post-measurement state when Eve observes the $l-$th outcome of her measurement $E_{1}$ is given by
$\rho_{A}^{l}=\frac{1}{\overline{P}(l|e=1)}\mathrm{Tr}_{E}\left[\left(\mathbbm{1}_{A}\otimes R_{l|1}\right)\bigotimes_{s=1}^{N}|\psi_{A_{s}E_{s}}\rangle\!\langle\psi_{A_{s}E_{s}}|\right].$ (154)
Then, substituting the states $|\psi_{A_{s}E_{s}}\rangle$ from Eq. (29) and the measurement elements $R_{l|1}$ from Eq. (30), we get that
$U_{A}\,\rho_{A}^{l}\,U_{A}^{\dagger}=\frac{1}{\overline{P}(l|e=1)}\mathrm{Tr}_{E}\left[\left(\mathbbm{1}_{A}\otimes(R^{\prime}_{l|1})^{*}\otimes\mathbbm{1}_{E^{\prime\prime}}\right)\bigotimes_{s=1}^{N}|\phi^{+}\rangle\!\langle\phi^{+}|_{A_{s}^{\prime}E_{s}^{\prime}}\otimes|\xi_{s}\rangle\!\langle\xi_{s}|_{A_{s}^{\prime\prime}E_{s}^{\prime\prime}}\right].$ (155)
Again using the identity $(\mathbbm{1}\otimes Q)|\phi^{+}\rangle=(Q^{T}\otimes\mathbbm{1})|\phi^{+}\rangle$, we obtain
$U_{A}\,\rho_{A}^{l}\,U_{A}^{\dagger}=\frac{1}{\overline{P}(l|e=1)}\mathrm{Tr}_{E}\left[\left(R^{\prime}_{l|1}\otimes\mathbbm{1}_{A^{\prime\prime}}\otimes\mathbbm{1}_{E}\right)\bigotimes_{s=1}^{N}|\phi^{+}\rangle\!\langle\phi^{+}|_{A_{s}^{\prime}E_{s}^{\prime}}\otimes|\xi_{s}\rangle\!\langle\xi_{s}|_{A_{s}^{\prime\prime}E_{s}^{\prime\prime}}\right],$ (156)
where we also used the fact that $\left((R^{\prime}_{l|1})^{*}\right)^{T}=R^{\prime}_{l|1}$. After tracing the $E$ subsystem we arrive at
$U_{A}\,\rho_{A}^{l}\,U_{A}^{\dagger}=\frac{1}{2^{N}\overline{P}(l|e=1)}R^{\prime}_{l|1}\otimes\tilde{\rho}_{A^{\prime\prime}},$ (157)
where $\tilde{\rho}_{A^{\prime\prime}}=\mathrm{Tr}_{E^{\prime\prime}}\left[\bigotimes_{s=1}^{N}|\xi_{s}\rangle\!\langle\xi_{s}|_{A_{s}^{\prime\prime}E_{s}^{\prime\prime}}\right]$. It is straightforward to observe now that $\overline{P}(l|e=1)=\mathrm{Tr}\left(R^{\prime}_{l|1}\right)/2^{N}$ giving us the desired result. ∎
Now, to certify an arbitrary pure state $|\psi\rangle\in(\mathbbm{C}^{2})^{\otimes N}$ among the external parties it is enough that the second Eve’s reference measurement $E_{1}^{\prime}$ is projective and one of its elements, say $R^{\prime}_{0|1}$, is a projection onto $|\psi\rangle$, $R^{\prime}_{0|1}=|\psi\rangle\!\langle\psi|$.
For mixed states the situation is a bit more complicated. Consider a mixed state acting on $\mathbbm{C}^{d}$:
$\rho=\sum_{k}p_{k}|\psi_{k}\rangle\!\langle\psi_{k}|.$ (158)
It follows that one can construct an extremal $3d-$outcome POVM in the Hilbert space $\mathbbm{C}^{2d}$ that can be performed by Eve on her share of the joint state $|\psi_{AE}\rangle$ to create $\rho$ at the external parties’ labs with the aid of post-processing. To construct this POVM, we first define two sets of $d$ mutually orthogonal vectors $\mathcal{B}_{1}=\{|\psi_{k}\rangle\}$ and $\mathcal{B}_{2}=\{|\phi_{k}\rangle\}$ such that $\langle\psi_{k}|\phi_{k^{\prime}}\rangle=0$ for any $k,k^{\prime}$. Then, one considers the pair of the states $\{\psi_{k},\phi_{k}\}$ to construct the trine POVM that acts on $\mathbbm{C}^{2}$ [cf. Ref. [33]] as
$M_{k,1}=p_{k}|\psi_{k}\rangle\!\langle\psi_{k}|,\qquad M_{k,2}=\frac{2-p_{k}}{2}|\tau_{k,2}\rangle\!\langle\tau_{k,2}|,\qquad M_{k,3}=\frac{2-p_{k}}{2}|\tau_{k,3}\rangle\!\langle\tau_{k,3}|,$ (159)
where for all $k$,
$|\tau_{k,2}\rangle=\sqrt{\frac{1-p_{k}}{2-p_{k}}}|\psi_{k}\rangle+\sqrt{\frac{1}{2-p_{k}}}|\phi_{k}\rangle,\qquad|\tau_{k,3}\rangle=-\sqrt{\frac{1-p_{k}}{2-p_{k}}}|\psi_{k}\rangle+\sqrt{\frac{1}{2-p_{k}}}|\phi_{k}\rangle.$ (160)
Notice that for any $k$,
$M_{k,0}+M_{k,1}+M_{k,2}=|\psi_{k}\rangle\!\langle\psi_{k}|+|\phi_{k}\rangle\!\langle\phi_{k}|,$ (161)
and therefore the three-element set $\{M_{k,l}\}_{l}$ can also be understood as a three-outcome rank-one POVM in $\mathbbm{C}^{2}$. Moreover, taking a sum over $k$ in Eq. (D8) one obtains that
$\sum_{k=0}^{d-1}(M_{k,0}+M_{k,1}+M_{k,2})=\sum_{k=0}^{d-1}(|\psi_{k}\rangle\!\langle\psi_{k}|+|\phi_{k}\rangle\!\langle\phi_{k}|)=\mathbbm{1}_{\mathbbm{C}^{2d}}$ (162)
which implies that $M_{k,l}$ form a valid $3d$-outcome rank-one POVM in $\mathbbm{C}^{2d}$ which we denote =$\mathcal{M}=\{M_{k,l}\}$.
Let us now demonstrate that $\mathcal{M}$ is an extremal POVM. To this end, we first notice that for any $k$ the elements $M_{k,l}$ $(l=0,1,2)$ are linearly independent and, moreover, the sets $\{M_{k,0},M_{k,1},M_{k,2}\}$ and $\{M_{k^{\prime},0},M_{k^{\prime},1},M_{k^{\prime},2}\}$ for any pair $k\neq k^{\prime}$ are orthogonal in the sense that any element from the first set is orthogonal to any element in the second set. This means that all $M_{k,l}$ forming our POVM are linearly independent which together with the fact that $M_{k,l}$ are rank-one implies that $\mathcal{M}$ is extremal [cf. Corollary 5 in Ref. [34]].
Let us now go back to our scheme of probabilistic self-testing of an arbitrary mixed state $\rho$ and observe that after performing $\mathcal{M}$ on her shares of the joint state, Eve obtains the outcomes $l\equiv(k,1)$ with probability $p(k,1|e=1)p_{l}/2^{N}$. Whenever she obtains the outcome $(k,1)$, the post-measurement state with the external parties is certified from Corrollary 1. Now, one can clearly see that as the average state with the external parties when Eve obtains $(k,1)$ for all $k$ is certified from Corrollary 1 to be
$U_{A}\,\rho_{A}\,U_{A}^{\dagger}=\frac{1}{\sum_{k}p(k,1)}\sum_{k}p(k,1)U_{A}\,\rho_{A}^{(k,1)}\,U_{A}^{\dagger}=\left(\sum_{k}M_{(k,1)}\right)\otimes\tilde{\rho}_{A^{\prime\prime}}=\rho_{A^{\prime}}\otimes\tilde{\rho}_{A^{\prime\prime}},$ (163)
or its complex conjugate, which is exactly as we promised. The above state occurs with a probability $1/2^{N}$.